### v1

# I2S-TFCKD: Intra-Inter Set Knowledge Distillation with Time-Frequency Calibration for Speech Enhancement

Abstract
In recent years, complexity compression of neural network (NN)-based speech enhancement (SE) models has gradually attracted the attention of researchers, especially in scenarios with limited hardware resources or strict latency requirements. The main difficulties and challenges lie in achieving a balance between complexity and performance according to the characteristics of the task. In this paper, we propose an intra-inter set knowledge distillation (KD) framework with time-frequency calibration (I2S-TFCKD) for SE. Different from previous distillation strategies for SE, the proposed framework fully utilizes the time-frequency differential information of speech while promoting global knowledge flow. Firstly, we propose a multi-layer interactive distillation based on dual-stream time-frequency cross-calibration, which calculates the teacher-student similarity calibration weights in the time and frequency domains respectively and performs cross-weighting, thus enabling refined allocation of distillation contributions across different layers according to speech characteristics. Secondly, we construct a collaborative distillation paradigm for intra-set and inter-set correlations. Within a correlated set, multi-layer teacher-student features are pairwise matched for calibrated distillation. Subsequently, we generate representative features from each correlated set through residual fusion to form the fused feature set that enables inter-set knowledge interaction. The proposed distillation strategy is applied to the dual-path dilated convolutional recurrent network (DPDCRN) that ranked first in the SE track of the L3DAS23 challenge. To evaluate the effectiveness of I2S-TFCKD, we conduct experiments on both single-channel and multi-channel SE datasets. Objective evaluations demonstrate that the proposed KD strategy consistently and effectively improves the performance of the low-complexity student model and outperforms other distillation schemes.
Index Terms:
speech enhancement, knowledge distillation, time-frequency calibration, intra-inter set correlations, residual fusion.
publicationid: pubid: $0$–$00$ 2025 IEEE
I Introduction
Speech enhancement (SE) aims to remove background noises and recover clean speech from noisy mixtures. As a perception front-end algorithm, its applications have already penetrated into diverse domains including such as intelligent terminals, industrial manufacturing, medical health, etc. [1]. Traditional SE methods based on statistical signal processing [2] estimate the noise spectrum from noisy speech for suppression, requiring low computational resources. However, these algorithms typically rely on stationarity assumptions, which fundamentally constrain their performance in handling non-stationary noises encountered in real-world scenarios [3]. With the recent emergence of data-driven deep learning approaches, SE models have begun leveraging massive training data to characterize both speech signals, noise components, and their nonlinear relationships. Capitalizing on the representational learning capabilities of deep neural networks (DNNs), deep learning-based SE methods have achieved superior noise suppression performances and have attained state-of-the-art results in international SE challenges [4], gradually becoming the mainstream SE solutions.
Recent SE models tend to adopt time-frequency (T-F) domain model architectures, decomposing speech signals into T-F units to capture local energy distribution and long-term correlations of speech. Typical T-F enhancement frameworks include the full-band and sub-band fusion network (FullSubNet) [5] that cascades a full-band model with a sub-band model to concurrently model local spectral patterns and long-distance cross-band dependencies. Frequency recurrent convolutional recurrent network (FRCRN) [6] applies frequency recursion to 3D convolutional feature maps along the frequency axis and utilizes a time-recurrent module to capture the temporal dynamics. These methods specifically tailor network architectures to the T-F characteristics of speech signals, comprehensively exploring the representation patterns of speech and noise to achieve noise suppression while preserving speech clarity. However, the superior performances of current NN-based SE methods often comes with substantial computational overhead. Top-ranked SE models in international challenges and public benchmarks typically have over 5M parameters and more than 10G floating-point operations (FLOPs) per second. Such excessive computational and memory requirements create deployment bottlenecks for edge devices, particularly in teleconferencing systems, Bluetooth headsets and hearing aids. Consequently, the exploration of SE models balancing performance with lightweight design has emerged as a research hotspot and trend.
Model compression is an effective solution, including pruning [7], quantization [8], and knowledge distillation (KD) [9]. Stamenovic et al. [10] specifically investigate pruning methodologies tailored for SE frameworks. Tan et al. [11] reduce the complexity of SE networks using three different techniques: sparse regularization, iterative pruning, and clustering-based quantization. However, pruning and quantization still heavily depend on the original model structure and hardware support. In contrast, KD transfers knowledge from a powerful teacher to a compact student, offering better generality and flexibility—thus becoming the technical path chosen in this paper. Actually, many researchers have already focused on distillation methods for SE models. Kim et al. [12] uses the teacher’s pseudo target to train the student model, enabling a zero-shot learning procedure. Thakker et al. [13] integrate KD with multi-task learning losses from speech recognition to achieve SE model compression. These studies primarily focus on minimizing output-level discrepancies between teachers and students while neglecting intermediate feature representations. Recently, Wan et al. [14] propose a hierarchical attention-based distillation framework for SE, exploring the utilization of intermediate representations. Similarly, Nathoo et al. [15] introduce a fine-grained similarity-preserving KD loss, which aims to match the student’s intra-activation Gram matrices to that of the teacher. Knowledge transfer of intermediate model representations has gradually attracted researchers’ attention, with the key challenge lying in the SE-specific distillation design.
In our previous work [16], we proposed a unified residual fusion probabilistic KD framework for SE, which leverages cross-layer supervision to guide information transfer between teacher and student while introducing a frame weighting probabilistic distillation loss to perceive key frame information. In this paper, we design an intra-inter set KD with time-frequency calibration (I2S-TFCKD) framework to further optimize the cross-layer supervision distillation paradigm. On the one hand, we introduce temporal and spectral domain cross-computation to derive multi-layer distillation calibration weights, leveraging time-frequency characteristics for targeted distillation optimization. On the other hand, unlike previous approaches that conduct cross-layer interactions within isolated correlation sets, we propose collaborative distillation across intra-set and inter-set correlations, effectively promoting knowledge flow throughout the model architecture. The backbone network for distillation employs our previously proposed dual-path dilated convolutional recurrent network (DPDCRN) model [17], which won the SE track of the L3DAS23 challenge. The main contributions of this work are as follows:
  • We propose multi-layer time-frequency cross-calibration distillation, which independently computes teacher-student similarity calibration weights along temporal and spectral dimensions for cross-weighting, enabling refined allocation of distillation contributions across different layers.
  • We design a collaborative distillation paradigm for intra-set and inter-set correlations. First, we partition the SE model structure into multiple correlation sets for intra-set multi-layer teacher-student matching. Subsequently, representative features from each correlated set are generated through residual recursive fusion, forming the fused feature set that enables inter-set knowledge interaction.
  • Extensive evaluations on public SE benchmarks demonstrate that our proposed I2S-TFCKD framework for SE surpasses existing distillation methods and enabling the student model to achieve competitive enhancement results with low computational complexity.
II Related Work
In this section, we provide an overview of related works on distillation methodologies for SE, particularly focusing on feature distillation-related research.
II-A Knowledge Distillation Strategy
KD originated from the pioneering work of Hinton et al. [18], where the student model learns from both ground-truth labels and soft labels provided by the teacher model. Building upon this, subsequent research focuses on exploring effective regularization methods to balance soft and hard labels. In recent years, feature distillation has gradually come into the spotlight, with numerous studies demonstrating that distilling intermediate feature representations exhibits stronger generalization ability and more promising results compared to soft-label distillation. Niu et al. [19] transform the feature maps of different levels into a frequency-enhanced feature space, which highlights the low-frequency features for both the teacher and student networks. Cross-image relational knowledge extraction is proposed in [20], leveraging pixel-to-pixel and pixel-to-region distillation to fully utilize the structural relationships between images. These methods primarily utilize single-layer features for knowledge transfer, while recent advances explore multi-layer and multi-stage feature distillation frameworks. A cross-layer knowledge review mechanism is investigated in [21], which integrates multi-layer teacher features to guide the student. Zhao et al. [22] developed a dense KD in all aspects manner, using the dynamic class-specific pseudo-labeling strategy to distill intermediate features and output logits collaboratively with more emphasis on semantic-invariant knowledge inheritance. Overall, feature distillation can transfer fine-grained knowledge such as edge, texture, and semantic features, garnering increasing research attention. Cross-layer feature distillation transcends traditional layer-wise correspondence constraints, allowing a student to simultaneously learn low-level detail features and high-level abstract representations from teachers. However, this paradigm also introduces challenges in representation alignment, making it a worthwhile direction for in-depth research.
II-B Distillation for Speech Enhancement
Most existing distillation works discussed above have been primarily applied to the computer vision and natural language processing domains. In the acoustic field, KD methods are initially adopted for classification tasks such as acoustic scene classification [23] and automatic speech recognition [24]. Subsequently, KD for regression tasks like SE has gradually emerged. Early distillation studies for SE models primarily focus on minimizing output-level discrepancies between teachers and students, exemplified by elite sub-band distillation [25]. These works aim to align student models with teacher patterns at the learning objective level but overlook the utilization of intermediate representations containing richer information. Feature distillation holds greater potential for SE models as their architectures and functionalities grow increasingly complex. A common challenge in representation-based distillation arises from dimension mismatch between teacher and student features. To address this, our previous work [26] first proposes a cross-layer similarity distillation framework for SE models, which aligns teacher-student features via similarity loss. Similarly, multiple studies have adopted multi-layer distillation strategies to compress SE models, including multi-view attention transfer distillation [27] and two-step fine-grained similarity-preserving KD [15]. It is worth noting that recent studies have begun to focus on leveraging the inherent characteristics of the SE task to improve distillation frameworks. Dynamic Frequency Adaptive Distillation [28] dynamically evaluates model outputs and adjusts distillation targets according to the requirements of different speech frequency bands. An independent attention transfer mechanism is designed for temporal and channel dimensions in [29] to facilitate rich knowledge transfer across different structured models. However, existing studies have not fully exploited the time-frequency characteristics of speech in designing feature distillation frameworks. Additionally, cross-layer distillation frameworks such as [15, 16] do not consider the global propagation of knowledge. To address these limitations, this work mainly includes two innovative improvements: the time-frequency calibration strategy for distillation and the intra-inter set global knowledge transfer mechanism.
III Methodology
III-A Teacher and Student Network Architecture
Figure 1: Backbone network architecture of the teacher model.
Figure 2: Overall architecture of the I2S-TFCKD framework. We present the detailed process of intra-inter set distillation. Fig. 2(a) shows the time-frequency cross-calibration knowledge transfer and the recursive feature fusion across different layers within a single correlated set. Fig. 2(b) demonstrates inter-set distillation achieved via the fused feature set among various correlated sets. Fig. 2(c) visualizes the process of generating fused features by utilizing the features of the current layer and the inherited recursive features.
In our previous work [16], we designed a residual fusion distillation framework for SE based on the DPDCRN architecture that won the SE track of L3DAS23 challenge. In this paper, we extend the multi-layer distillation within correlated sets to intra-inter set global knowledge transfer, while implementing time-frequency cross-assignment weighting in the distillation loss. The backbone architecture remains DPDCRN due to its facility for structural compression and scalability to both single-channel and multi-channel SE tasks. The backbone architecture of the teacher model is presented in Fig. 1. The main structure of DPDCRN consists of a convolutional encoder, a convolutional decoder, and an intermediate frequency-time (F-T) processing module. The encoder contains two convolutional layers and four dilated convolutional blocks, while the decoder contains two deconvolutional layers and four dilated convolutional blocks. Each frequency-time processing module contains dual branches for frequency and temporal axes, with each branch comprising a causal multi-head self-attention (MHSA) module and a feedforward network based on gated recurrent units (GRUs). For each convolutional layer, the student has half the number of channels of the teacher. For the F-T processing modules, the number of hidden units in the student is half that of the teacher. Additionally, the teacher model has four cascaded F-T modules, while the student only uses one. Overall, the student contains merely 17% of the teacher’s parameters.
III-B Intra-inter Set Knowledge Transfer Architecture
According to the encoder-decoder framework of current mainstream SE models, we partition all layers of teacher-student networks into three correlated sets corresponding to the encoder, decoder, and intermediate F-T processing modules, respectively, to achieve hierarchical knowledge transfer. In previous work [16], we utilized residual fusion to implement cross-layer knowledge transfer within each correlated set and confirmed its advantages over conventional layer-wise distillation. However, residual fusion distillation introduces redundant replication operations when the number of modules in teacher and student networks differs [16], and cross-set information interaction remains unexplored. In this paper, we propose a unified cross-layer distillation framework that jointly addresses intra-set and inter-set knowledge transfer. For intra-set operations, single-layer student features are aligned with all corresponding teacher representations through weighted matching within the correlated set. Regarding inter-set transfer, we first generate representative fusion features for each set via residual fusion, then perform multi-layer knowledge transfer on the fused feature set to facilitate cross-hierarchical circulation of intermediate representations. The overall architecture is shown in Fig. 2.
III-C Time-frequency Cross-calibrated Multi-layer Matching Distillation
The candidate correlation set of the teacher and student is denoted as ${\cal P}=\left\{{\left({{l_{s}},{l_{t}}}\right)\left|{\forall{l_{s}}\in\left[{1,\ldots,{L_{s}}}\right]}\right.,{l_{t}}\in\left[{1,\ldots,{L_{t}}}\right]}\right\}$, where ${{l_{s}}}$ and ${{l_{t}}}$ represent the corresponding layers of the candidate student and teacher, respectively. In the feature distillation paradigm, the feature maps of candidate teacher-student pairs are first transformed into specific embedding representations, and then, the distance between teacher-student embeddings is calculated as the distillation loss ${{\cal L}_{KD}}$:
${{\cal L}_{KD}}=\sum\limits_{\left({{l_{s}},{l_{t}}}\right)\in{\cal P}}{{\cal D}\left({Tran{s^{s}}\left({{{\bf{F}}_{{l_{s}}}}}\right),Tran{s^{t}}\left({{{\bf{F}}_{{l_{t}}}}}\right)}\right)},$ (1)
where ${{\bf{F}}_{{l_{s}}}}$ and ${{\bf{F}}_{{l_{t}}}}$ denote the hidden layer feature representations of the student and teacher respectively, $Tran{s^{s}}$ and $Tran{s^{t}}$ are used to transform teacher-student feature pairs into specific embedding representations for alignment. Conventional layer-wise distillation imposes strict constraints on structural and layer alignment between teacher and student models, requiring redundant layer information to be discarded in cases of mismatch. In contrast, multi-layer distillation that matches single-layer student features with multi-layer teacher features, is not constrained by the difference in the number of layers.
Different layer structures in DNNs exhibit varying capabilities to learn feature representations. As the depth of layers increases, the intermediate representation information of the model gradually becomes more abstract. Although multi-layer feature matching enables student models to perceive a broader range of teacher hidden knowledge, indiscriminate alignment may cause semantic redundancy that degrades learning efficiency and performance. To maximize absorption of valuable information, student layers should preferentially align with teacher layers exhibiting high semantic relevance. Inspired by the observation that the proximity of pairwise similarity matrices can be regarded as a good measurement of the inherent semantic similarity [30], we use semantic-aware similarity for distillation calibration. Specifically, we introduce a calibration weight ${\alpha_{\left({{l_{s}},{l_{t}}}\right)}}$ to dynamically adjust the contribution of candidate teacher-student layer pairs based on their semantic congruence, and constrain the corresponding weights through a softmax function: $\sum\limits_{{l_{t}}=1}^{{L_{t}}}{{\alpha_{\left({{l_{s}},{l_{t}}}\right)}}}=1,\forall{l_{s}}\in\left[{1,\ldots,{L_{s}}}\right]$.
For SE models, time-frequency interaction analysis enables simultaneous capture of local spectral variations and temporal dynamics in speech signals, which is expected to facilitate more refined alignment of student and correlated teachers. Therefore, we compute calibration weights for temporal and spectral flows independently and perform cross-weighting. Given a batch of feature map input ${{\bf{F}}_{s/t}}\in{{\mathop{\rm R}\nolimits}^{B\times C\times T\times D}}$, where $B$ is the batch size, $C$ is the number of output channels, $T$ is the number of speech frames, $D$ is the feature dimension size, and $s/t$ denotes the student or teacher, we first calculate the similarity maps in the time-flow and frequency-flow respectively, as shown in Fig. 3. For the time-domain flow, the feature map is split into multiple instances along the batch dimension, then flattened along the channel and feature dimensions. ${l^{2}}$ Normalization is applied to the merged feature dimension to obtain the temporal transformed features ${\bf{M}}_{s/t}^{T}\in{{\mathop{\rm R}\nolimits}^{B\times T\times\left({C\times D}\right)}}$. Finally, the time-flow mapping matrix ${\bf{P}}_{s/t}^{T}\in{{\mathop{\rm R}\nolimits}^{B\times T\times T}}$ is computed using cosine similarity measurement:
${\bf{P}}_{s/t}^{T}=\frac{1}{2}\left({\frac{{{{\left({{\bf{M}}_{s/t}^{T}}\right)}^{\mathop{\rm T}\nolimits}}\cdot{\bf{M}}_{s/t}^{T}}}{{\left\|{{\bf{M}}_{s/t}^{T}}\right\|_{2}^{2}}}+1}\right).$ (2)
For the frequency-domain flow, we partition the feature maps along the time dimension into multiple instances to maintain frame independence. Then, the frequency transformed features ${\bf{M}}_{s/t}^{F}\in{{\mathop{\rm R}\nolimits}^{T\times B\times\left({C\times D}\right)}}$ are obtained by flattening and normalization. Finally, the frequency-flow mapping matrix ${\bf{P}}_{s/t}^{F}\in{{\mathop{\rm R}\nolimits}^{T\times B\times B}}$ is also calculated by cosine similarity.
Subsequently, we employ trainable weight matrices to project the teacher-student self-similarity mapping matrices from both the temporal and spectral domains into query and key subspaces, mitigating the effects of noise and sparsity:
${\bf{Q}}_{{l_{s}}}^{T/F}=EM{B_{Q}}\left({{\bf{P}}_{{l_{s}}}^{T/F}}\right){\rm{,}}$ (3)
${\bf{K}}_{{l_{t}}}^{T/F}=EM{B_{K}}\left({{\bf{P}}_{{l_{t}}}^{T/F}}\right),$
where $EM{B_{Q}}\left(\cdot\right)$ and $EM{B_{K}}\left(\cdot\right)$ consist of the fully connected layer, ReLU activation, and layer normalization. The time-domain and frequency-domain flows adopt independent trainable weight parameters. The similarity calibration coefficients for the time and frequency domains are computed through domain-specific query-key interactions:
$\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{T/F}={e^{{{\left({{\bf{Q}}_{{l_{s}}}^{T/F}}\right)}^{\rm{T}}}{\bf{K}}_{{l_{t}}}^{T/F}}}\cdot{\left({\sum\limits_{{l_{t}}\in{\cal P}}{{e^{{{\left({{\bf{Q}}_{{l_{s}}}^{T/F}}\right)}^{\rm{T}}}{\bf{K}}_{{l_{t}}}^{T/F}}}}}\right)^{-1}}.$ (4)
Finally, cross-layer matching is performed for multiple teacher-student feature pairs within the current correlated set, where similarity metrics are computed through both time and frequency flows and distillation contributions are allocated via $T/F$ calibration coefficients. The multi-layer time-frequency calibration knowledge distillation (TFCKD) loss is calculated as follows:
${\cal L}_{{\rm{TFCKD}}}$ $=\sum\limits_{{l_{s}}=1}^{{L_{s}}}\sum\limits_{{l_{t}}=1}^{{L_{t}}}\left(\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{T}\left({{\bf{P}}_{{l_{t}}}^{T}-{\bf{P}}_{{l_{s}}}^{T}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{T}}}{{{\bf{P}}_{{l_{s}}}^{T}}}}\right)\right.$ (5)
$\left.+\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{F}\left({{\bf{P}}_{{l_{t}}}^{F}-{\bf{P}}_{{l_{s}}}^{F}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{F}}}{{{\bf{P}}_{{l_{s}}}^{F}}}}\right)\right),$
where ${L_{s}}$ and ${L_{t}}$ are the total number of layers of the student and teacher in the current correlated set. This distillation loss serves as the knowledge transfer mechanism for both intra-set and inter-set teacher-student features.
Figure 3: Similarity mapping of time and frequency flows. The self-similarity matrices for distillation are calculated in two flows: the time domain and the frequency domain. The distribution of features along the time axis will affect the similarity computation in the time flow, whereas the frequency domain processing operates frame-wise.
III-D Residual Fusion Features for Correlated Sets
We partition the teacher-student correlated sets based on the functionality of structural blocks in the SE model. Distillation within each set enables the student to focus on the teacher’s representational information in the corresponding functional region. Meanwhile, information interaction between correlated sets is expected to facilitate global knowledge circulation. As illustrated in Fig. 2(b), we first generate representative features for each correlated set through residual fusion to construct a teacher-student fused feature set. Then, the multi-layer distillation with time-frequency calibration is performed within this set. Fused features are generated via binary fusion of the current layer’s features and recursive features are inherited from the previous layer, and are continuously propagated forward. The process of single-round residual fusion is shown in Fig. 2(c).
The feature representation at the current $j$-th layer is denoted as ${{\bf{F}}^{j}}$, and the recursive features inherited from the previous layer are ${{\bf{R}}^{j-1}}$. First, the representation dimensions are aligned via 2D convolution and up-sampling operations, respectively:
${{{\bf{\tilde{R}}}}^{j-1}}={\rm{2DConv}}\left({{{\bf{R}}^{j-1}}}\right),$ (6)
${{{\bf{\tilde{F}}}}^{j}}={\rm{UpSampling}}\left({{{\bf{F}}^{j}}}\right),$
where the up sampling operation is used to transform the feature dimensions, while the 2D convolution is employed to adjust the channel dimensions. Subsequently, the transformed layer features and the recursive features are concatenated along the channel dimension, and an attention vector is generated via 1D convolution:
${{\bf{A}}^{j}}=\sigma\left({1{\rm{DConv}}\left({{\rm{Concat}}\left({{{{\bf{\tilde{F}}}}^{j}},{{{\bf{\tilde{R}}}}^{j-1}}}\right)}\right)}\right),$ (7)
where $\sigma$ represents the sigmoid activation function. The attention vector has two channels corresponding to the retention weight coefficients ${\bf{A}}_{\bf{F}}^{j}$ and ${\bf{A}}_{\bf{R}}^{j}$ for the layer features and recursive features, respectively. The recursive feature ${{\bf{R}}^{j}}$ of the current layer is generated by the weighted summation of the two branches:
${{\bf{R}}^{j}}={\bf{A}}_{\bf{R}}^{j}\cdot{{\bf{\tilde{R}}}^{j-1}}+{\bf{A}}_{\bf{F}}^{j}\cdot{{\bf{\tilde{F}}}^{j}},$ (8)
where ${{\bf{R}}^{j}}$ will continue to participate in recursive computation as the fusion input for the next layer, while the fused output of the module is generated through 2D convolution:
${{\bf{T}}^{j}}={\rm{2DConv}}\left({{{\bf{R}}^{j}}}\right).$ (9)
For each correlated set, we select the last recursion output as the representative fusion feature of the current set, which is then integrated to form a fused feature set. We believe that due to the symmetry of the encoder-decoder architecture in SE models, higher-level features from the middle layers exhibit stronger capability to learn effective information from lower-level features near the input/output ends. Therefore, in terms of residual fusion direction, the encoder and intermediate F-T processing modules follow a forward order starting from the first layer, while the decoder adopts a reverse order integrating from the final layer towards the middle. Corresponding representative features are generated for each correlated set in both the teacher and student models following the above process, and integrated into the teacher-student fused feature set for inter-set distillation.
III-E Training Procedure
In this section, we elaborate on the entire training procedure for intra-inter set distillation. The overall workflow is outlined in Algorithm 1. First, we pretrain the teacher network using the multi-resolution short-time Fourier transform (MRSTFT) loss as the backbone loss. The teacher is then frozen to serve as a guide for the student, and KD proceeds concurrently with the student’s training. Intra-set distillation is applied to the encoder, decoder, and F-T processing modules. Within each correlated set, besides calculating the intra-set time-frequency calibration distillation loss ${{\cal L}_{Intra}}$, we generate representative teacher-student fused features for the current set through layer-wise recursive computation. Inter-set distillation loss ${{\cal L}_{Inter}}$ is derived via cross-layer interactions within the fused feature set. The total loss for training the student model is formulated as:
${{\cal L}_{stu}}$ $={{\cal L}_{MRSTFT}}+\sum\limits_{k=1}^{q}{\sum\limits_{i=1}^{m}{\sum\limits_{j=1}^{n}{{{\cal L}_{Intra}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)}}}$ (10)
$+\sum\limits_{i=1}^{q}{\sum\limits_{j=1}^{q}{{{\cal L}_{Inter}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)}},$
where $q$ is the total number of correlated sets (set to 3 in this paper), $m$ is the number of layers contained in the student correlated set, and $n$ is the number of layers contained in the teacher correlated set.
IV Experimental Settings
IV-A Dataset Setup
Input: Student and teacher correlated sets $\left\{{C_{s}^{k}}\right\}_{k=1}^{q}$ and $\left\{{C_{t}^{k}}\right\}_{k=1}^{q}$.
Output: Intra-set and inter-set distillation loss ${{\cal L}_{Intra}}$ and ${{\cal L}_{Inter}}$.
/* Iterate through all correlated sets. */
1 for $k\leftarrow 1$ to $q$ do
2       Sample intra-set student representations $\left({{\bf{F}}_{s}^{\rm{1}},{\bf{F}}_{s}^{2}\ldots,{\bf{F}}_{s}^{m}}\right)$ from $C_{s}^{k}$;
3       Sample intra-set teacher representations $\left({{\bf{F}}_{t}^{\rm{1}},{\bf{F}}_{t}^{2}\ldots,{\bf{F}}_{t}^{n}}\right)$ from $C_{t}^{k}$;
4       ${{\cal L}_{Intra}}=0$;
       /* Iterate through all teacher-student layers in the current set. */
5       for $i\leftarrow 1$ to $m$ do
6             for $j\leftarrow 1$ to $n$ do
7                   ${{\cal L}_{Intra}}+={\rm{TF\_Calibration}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)$;
8                   if $i==m$ then // Ensure the uniqueness.
9                         if $j==1$ then
10                               ${\bf{R}}_{t}^{j}={\bf{F}}_{t}^{j}$;
11
12                        else
13                               ${\bf{T}}_{t}^{j},{\bf{R}}_{t}^{j}={\rm{Fusion}}\left({{\bf{F}}_{t}^{j},{\bf{R}}_{t}^{j-1}}\right)$;
14
15                         end if
16
17                   end if
18
19             end for
20            if $i==1$ then
21                   ${\bf{R}}_{s}^{j}={\bf{F}}_{s}^{j}$;
22
23            else
24                   ${\bf{T}}_{s}^{j},{\bf{R}}_{s}^{j}={\rm{Fusion}}\left({{\bf{F}}_{s}^{j},{\bf{R}}_{s}^{j-1}}\right)$;
25
26             end if
27
28       end for
29      ${\bf{u}}_{s}^{k}={\bf{T}}_{s}^{m}$;
       // The final output is chosen as the representative fused feature.
30       ${\bf{u}}_{t}^{k}={\bf{T}}_{t}^{n}$;
31
32 end for
[…]
/* Iterate through fused features. */
34 for $i\leftarrow 1$ to $q$ do
35       for $j\leftarrow 1$ to $q$ do
36             ${{\cal L}_{Inter}}+={\rm{TF\_Calibration}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)$;
37
38       end for
39
40 end for
Algorithm 1 Intra-inter Set Distillation Processing
The proposed distillation framework is trained and evaluated on two widely used public SE datasets: the single-channel deep noise suppression (DNS) challenge dataset [^1] The dataset is available at https://github.com/microsoft/DNS-Challenge. and the multi-channel L3DAS23 challenge dataset [^2] The dataset is available at https://www.l3das.com/icassp2023.. First, we conduct ablation studies on the DNS dataset to verify the effectiveness of the proposed distillation components. Subsequently, we compare our method with state-of-the-art (SOTA) models on the DNS official test set. Finally, the effectiveness of the proposed framework is further verified on the multi-channel L3DAS23 dataset.
The DNS dataset contains approximately 500 hours of clean speech and 180 hours of noise clips. We randomly partition the corpus into 60000 and 100 utterances for training and validation, respectively. Noisy speech samples are generated by mixing clean speech with noise at random signal-to-noise ratios (SNRs) ranging from -5dB to 15dB, utilizing 100 hours of speech data in total. For ablation studies, we construct a multi-SNR test set, where 100 speech clips that do not overlap with the training and validation sets are mixed with random noises at three SNR levels (-5dB, 0dB, and 5dB) to form 300 noisy-clean pairs. Additionally, the official non-reverb test set is used for comparisons of objective speech evaluation metrics between different algorithms.
The proposed distillation method’s performance on the multi-channel SE task is tested in the 3D SE track of the L3DAS23 challenge. The speech data contains over 4,000 virtual 3D audio clips, each lasting up to 12 seconds. The noise set includes 12 categories of transient noise and 4 categories of continuous noise. The data are released as two first-order ambisonics recordings (each with 4 channels). We divide the training data that contains a total of 80 hours of noisy-clean speech pairs into training and validation sets at a ratio of 75:1. Since the official blind test set is not fully open-sourced, we compare various algorithms on the 7-hour development test set provided by the challenge.
IV-B Implementation Details
TABLE I: Baseline model hyperparameter settings.
Layer Teacher Student
Parameters Input Size Output Size Parameters Input Size Output Size
Conv2d (1, 3), (1, 2), 128 ($C$, $T$, $F$) (128, $T$, $F/2$) (1, 3), (1, 2), 64 ($C$, $T$, $F$) (64, $T$, $F/2$)
Conv2d (1, 3), (1, 2), 128 (128, $T$, $F/2$) (128, $T$, $F/4$) (1, 3), (1, 2), 64 (64, $T$, $F/2$) (64, $T$, $F/4$)
Dilated Block (4$\times$×Conv2d) (2, 3), (1, 1), 128 $p$: [1, 2, 4, 8] (128, $T$, $F/4$) (128, $T$, $F/4$) $(2,3)$, $(1,1)$, 64 $p$: [1, 2, 4, 8] (64, $T$, $F/4$) (64, $T$, $F/4$)
Frequency-Time Block (128 GRU×$\times$2)$\times$4 (128, $T$, $F/4$) (128, $T$, $F/4$) 64 GRU×$\times$2 (64, $T$, $F/4$) (64, $T$, $F/4$)
Dilated Block (4$\times$×Conv2d) (2, 3), (1, 1), 128 $p$: [1, 2, 4, 8] (128, $T$, $F/4$) (128, $T$, $F/4$) $(2,3)$, $(1,1)$, 64 $p$: [1, 2, 4, 8] (64, $T$, $F/4$) (64, $T$, $F/4$)
DeConv2d (1, 3), (1, 2), 128 (128, $T$, $F/4$) (128, $T$, $F/2$) (1, 3), (1, 2), 64 (64, $T$, $F/4$) (64, $T$, $F/2$)
DeConv2d (1, 3), (1, 2), 2 (128, $T$, $F/2$) (2, $T$, $F$) (1, 3), (1, 2), 2 (64, $T$, $F/2$) (2, $T$, $F$)
TABLE II: Distillation module hyperparameter settings.
Similarity Embedding Mapping
Layer Time Flow Frequency Flow
Parameters Input Size Output Size Parameters Input Size Output Size
Linear $T*factor$ units ($B$, $T$, $T$) ($B$, $T$, $T*factor$) $B*factor$ units ($T$, $B$, $B$) ($T$, $B$, $B*factor$)
ReLU - ($B$, $T$, $T*factor$) ($B$, $T$, $T*factor$) - ($T$, $B$, $B*factor$) ($T$, $B$, $B*factor$)
Linear $T$ units ($B$, $T$, $T*factor$) ($B$, $T$, $T$) $B$ units ($T$, $B$, $B*factor$) ($T$, $B$, $B$)
${l^{2}}$orm - ($B$, $T$, $T$) ($B$, $T$, $T$) - ($T$, $B$, $B$) ($T$, $B$, $B$)
Residual Fusion
Layer Teacher Student
Parameters Input Size Output Size Parameters Input Size Output Size
Conv2d (3, 3), (1, 1), 128 ($C_{t}^{j}$, $T$, $F_{t}^{j}$) (128, $T$, $F_{t}^{j}$) (3, 3), (1, 1), 64 ($C_{s}^{j}$, $T$, $F_{s}^{j}$) (64, $T$, $F_{s}^{j}$)
Up Sampling - (128, $T$, $F_{t}^{j-1}$) (128, $T$, $F_{t}^{j}$) - (128, $T$, $F_{s}^{j-1}$) (64, $T$, $F_{s}^{j}$)
Concatenate - 2$\times$(128, $T$, $F_{t}^{j}$) (256, $T$, $F_{t}^{j}$) - 2$\times$(64, $T$, $F_{s}^{j}$) (128, $T$, $F_{s}^{j}$)
Conv1d 1, 1, 2 (256, $T$, $F_{t}^{j}$) (2, $T$, $F_{t}^{j}$) 1, 1, 2 (128, $T$, $F_{s}^{j}$) (2, $T$, $F_{s}^{j}$)
Sigmoid - (2, $T$, $F_{t}^{j}$) 2$\times$(1, $T$, $F_{t}^{j}$) - (2, $T$, $F_{s}^{j}$) 2$\times$(1, $T$, $F_{s}^{j}$)
Conv2d (3, 3), (1, 1), $C_{t}^{j}$ (128, $T$, $F_{t}^{j}$) ($C_{t}^{j}$, $T$, $F_{t}^{j}$) (3, 3), (1, 2), $C_{s}^{j}$ (64, $T$, $F_{s}^{j}$) ($C_{s}^{j}$, $T$, $F_{s}^{j}$)
In this paper, we use the DPDCRN model as our baseline, which takes complex spectrograms as input and outputs complex ratio masks. For the single-channel DNS dataset, the input channels are set to 2, while for the multi-channel L3DAS23 dataset, the input channels are set to 16. The hyperparameters of the teacher and student models are shown in Table I, where $C$ represents the number of input channels, $T$ is the number of speech frames, and $F$ is the frequency dimension of speech features. For the encoder and decoder parts, the number of channels in each convolutional layer of the teacher is twice that of the student. In Table I, $p$ represents the dilation factor, and both the teacher and student models use four types of dilation receptive fields. For the intermediate part, the teacher model employs four F-T modules, while the student model uses only one F-T module, with the number of GRU hidden units also being half those of the teacher model’s.
The hyperparameter settings for the trainable layers in the proposed distillation components are shown in Table II. The trainable components contain the computation of query and key embedding matrices in time-frequency calibration, as well as the generation process of recursively fused features for each correlated set. In the time-frequency calibration, the similarity embedding mappings are computed separately from two flow directions, with each flow containing two fully-connected layers for dimension transformation. The dimension transformation coefficient $factor$ in Table II is set to four in this paper. Residual fusion integrates the current layer features with the recursive features inherited from the previous layer via trainable parameters. In Table II, $C_{t/s}^{j}$ represents the number of teacher-student channels in the current layer, $F_{t/s}^{j}$ denotes the feature dimension of teacher-student features in the current layer, and $F_{t/s}^{j-1}$ denotes the feature dimension in the previous layer. The number of channels for recursive features of the teacher and student are set to 128 and 64, respectively. Notably, the trainable parameters of the distillation components are only involved in the training of the student model and do not impose additional burden during inference.
All speech signals are sampled at 16 kHz and segmented into 2.5-second chunks. For all models, the window length is set to 32 ms with a frame shift of 16 ms, and the FFT size is 512 points. We employ the MRSTFT loss [31] as the backbone loss. The models are implemented using PyTorch. The training configuration includes a learning rate of 0.0006 with the Adam optimizer, a batch size of 8, and a total training duration of 20 epochs. We provide source codes to ensure reproducibility [^3] The codes are available at https://github.com/JMCheng-SEU/I2S-TFCKD-SE..
IV-C Comparative Methods
We conduct multiple comparative experiments to comprehensively evaluate the proposed distillation framework for SE. For the comparative experiments on KD methods, we consider the following seven frameworks: MSE [32] that directly narrows the distance between teacher-student outputs; FitNet [33] that performs knowledge transfer between teacher and student through feature distillation of corresponding layers; SemCKD [30], a cross-layer semantic supervision distillation with learned attention distributions; CLSKD [26] that leverages cross-layer similarity between the teacher and the student for knowledge transfer; UCLFWPKD [16], a frame-weighting residual probabilistic KD framework; ABC-KD [14] that enables adaptive learning of compressed multi-layer knowledge through layer-wise attention mechanisms; Two-Step KD [15] that adopts a two-stage distillation framework, first pre-training the student using only KD criteria to match teacher activation patterns, followed by further optimization through supervised training.
Furthermore, we compare multiple state-of-the-art algorithms on the two benchmarks, respectively, to verify the competitiveness of the low-complexity student model after distillation. For the DNS benchmark, we select eight SE models for comparison. Specifically, RNNoise [34] uses gated recurrent networks (GRUs) to compute speech band gains. NSNet [35] introduces two SE loss functions based on mean squared error. DTLN [36] integrates analysis-synthesis methods through two cascaded networks. DCCRN [37] reconstructs the convolutional and recurrent network structure using complex-valued computations. FullSubNet+ [38] is an improved version of FullSubNet [5], strengthening frequency-band discriminability through multi-scale convolutions and channel attention. FRCRN [6] introduces a frequency-recurrent module applied to 3D convolutional feature maps along the frequency axis. CTS-Net [39] introduces a two-stage SE network, estimating magnitude spectral information in the first stage and suppressing residual noise while correcting phase information in the second stage. PrimeK-Net [40] utilizes the group prime-kernel feed-forward network to efficiently process time-frequency domain information.
For the L3DAS23 benchmark, we choose the following five comparative models. Neural Beamforming [41] serves as the baseline of the challenge, combining a beamforming framework designed for distributed microphones with deep neural networks. EaBNet [42] designs two modules: an embedding module that learns a 3D embedding tensor to represent time-frequency information and a beamforming module that derives beamforming weights to achieve filter-and-sum operations. CCA Speech [43] is a U-Net-based streaming attention framework that ranked third in the L3DAS23 challenge. DeFT-AN [44] incorporates three different types of blocks for aggregating information in the spatial, spectral, and temporal dimensions. Spatial Net [45] primarily consists of interleaved narrow-band and cross-band blocks to respectively exploit narrow-band and cross-band spatial information.
IV-D Evaluation Metrics
In this paper, we use six objective speech quality evaluation metrics, including the perceptual evaluation of speech quality (PESQ) [46], short-time objective intelligibility (STOI) [47], scale-invariant signal-to-noise ratio (SI-SNR) [48], and three mean opinion ratings on quality scales [49]: signal distortion (CSIG), the intrusiveness of background noise (CBAK), and overall quality (COVL). For all these metrics, higher scores indicate better performance. Additionally, on the L3DAS23 dataset, we use the official recommended word error rate (WER) computed via Wav2Vec [50] to evaluate SE effects for speech recognition purposes, where lower values indicate better performance.
V Experimental Results and Discussion
V-A Ablation Study
We conduct ablation experiments for the distillation components on both the validation set and the multi-SNR test set to verify the effectiveness of the proposed modules step-by-step. First, we set the M1 method as the basic form of feature distillation, adopting the teacher-student layer-wise distillation framework of FitNet. Then, method M2 is built upon the cross-layer attention-calibration distillation framework proposed by SemCKD, leveraging multi-layer teacher feature information to guide the student model. Method M3 expands on M2’s framework by introducing the multi-layer matching distillation with time-frequency calibration proposed in Section III-C of this paper. Finally, the M4 method (i. e., the proposed I2S-TFCKD framework) adds correlated set residual fusion to M3, and performs intra-inter set multi-layer distillation.
Figure 4: Training curve trends on the DNS validation set.
Figure 5: Frame-level heatmap distribution of time-frequency alignment weights.
V-A1 Analysis of Training Curves Trend
Fig. 4(a) shows the trend of each method’s backbone loss with the training epochs on the validation set. It can be observed that the feature distillation M1 enables the student’s backbone loss to decrease rapidly in the early training stages, but this initial advantage does not persist until model convergence. In contrast, method M2 achieves a lower converged backbone loss, indicating that cross-layer interactions provide additional knowledge to sustain the student’s learning. After introducing the time-frequency cross-calibration distillation loss in the M3 method, the variation of backbone loss with training epochs becomes more uniform and converges better, indicating that assigning multi-layer distillation weights based on time-frequency calibration leads to more stable knowledge transfer. On this basis, the M4 method introducing residual fusion distillation between correlated sets leads to lower convergence of the backbone loss. This demonstrates that intra-set and inter-set distillation complement each other, expanding the depth of knowledge perceived by the student model. Fig. 4(b) shows the variation of the speech metric PESQ with training epochs. The general performance order of different distillation methods is M1<M2<M3<M4, consistent with that in Fig. 4(a).
V-A2 Heatmap Visualization Analysis of Time-frequency Calibration Weights
We extract the trained time-frequency cross-calibration weights for a speech segment and compute the frame-level average energy along the feature dimension. Heatmap analysis is performed along the temporal and frequency domains, respectively, integrated with spectrograms of noisy and clean speech, as illustrated in Fig. 5. It can be observed that the temporal calibration weights allocate higher weights to frame regions with speech activity, whereas the weights computed along the frequency domain show no obvious focusing bias. This indicates that the temporal calibration weights can capture contribution differences across different speech frames, whereas frequency branch focuses on analyzing distillation contributions between different layers within the current frame.
V-A3 Analysis of Evaluation Metrics
TABLE III: Speech evaluation metrics for ablation comparisons of distillation components under different SNRs.
Methods Feature distillation Cross-layer distillation T-F calibration Intra-inter interactions Metric
PESQ CSIG CBAK COVL STOI SISNR
-5dB SNR
Noisy - - - - 1.289 2.066 1.828 1.612 0.751 1.541
Base model:
DPDCRN-T - - - - 2.249 3.720 3.058 2.990 0.935 11.744
DPDCRN-S - - - - 1.927 3.325 2.281 2.609 0.864 10.653
KD methods:
-M1 - - - 1.966 3.377 2.840 2.657 0.870 0.844
-M2 - - - 2.003 3.437 2.884 2.708 0.874 11.020
-M3 - - 2.057 3.500 2.894 2.767 0.880 11.107
-M4 - 2.106 3.567 2.950 2.830 0.886 11.253
0dB SNR
Noisy - - - - 1.380 2.407 2.108 1.844 0.825 6.457
Base model:
DPDCRN-T - - - - 2.644 4.088 3.426 3.393 0.940 14.717
DPDCRN-S - - - - 2.261 3.708 3.177 2.990 0.917 13.936
KD methods:
-M1 - - - 2.329 3.772 3.224 3.059 0.920 14.106
-M2 - - - 2.368 3.826 3.262 3.107 0.924 14.237
-M3 - - 2.426 3.904 3.274 3.174 0.928 14.270
-M4 - 2.491 3.954 3.333 3.238 0.931 14.461
5dB SNR
Noisy - - - - 1.546 2.796 2.475 2.140 0.886 11.343
Base model:
DPDCRN-T - - - - 2.992 4.400 3.760 3.744 0.964 17.327
DPDCRN-S - - - - 2.603 4.070 3.530 3.359 0.949 16.875
KD methods:
-M1 - - - 2.700 4.147 3.591 3.452 0.952 17.041
-M2 - - - 2.724 4.180 3.617 3.483 0.953 17.149
-M3 - - 2.788 4.244 3.633 3.546 0.956 17.157
-M4 - 2.850 4.285 3.686 3.606 0.958 17.356
Average SNR
Noisy - - - - 1.405 2.423 2.137 1.865 0.821 6.447
Base model:
DPDCRN-T - - - - 2.628 4.069 3.415 3.376 0.946 14.596
DPDCRN-S - - - - 2.264 3.701 3.173 2.986 0.910 13.821
KD methods:
-M1 - - - 2.332 ($\ast$)1 3.765 ($\ast$) 3.218 ($\ast$) 3.056 ($\ast$) 0.914 ($\ast$) 13.997 ($\ast$)
[…]
  • 1
[…]
Table III presents the comparison of objective speech metrics across different distillation strategies on the multi-SNR test set, and statistical significance analysis is conducted in the average metric comparison under three SNRs. We use the significance level $p$ to measure the statistical improvement of each distillation method compared to the student model. From the results, all distillation methods improve the student’s objective metrics across the three SNR levels, narrowing the gap between the teacher and student. Among them, M1 that performs layer-wise distillation, underperforms other methods, showing no significant improvement ($p>0.1$) over the student model in any of the six metrics. In contrast, M2 that extends single-layer distillation to attention-based multi-layer distillation, outperforms M1 across all metrics and achieves secondarily significant improvements ($0.10.05<p<$ 0.1) in CSIG and STOI compared to the student model. This indicates that cross-layer interactions provide additional representational knowledge. Building on this, M3 that introduces time-frequency cross-calibration, achieves further metric advantages, with statistically significant improvements ($p<0.05$) in CSIG, COVL, and STOI compared to the student model. This demonstrates that cross-calibration captures discriminative information across frequency bands and temporal segments at different layers, enabling more refined teacher-student layer matching. M4 that introduces residual fusion for inter-set distillation, achieves the best metric results among all strategies, with statistically significant improvements ($p<0.05$) in five metrics under average SNR compared to the student, except for SISNR. This indicates that collaborative distillation within and between correlated sets effectively promotes global knowledge circulation, significantly improving the enhancement effect of the student model.
V-B Comparative Experiments on the DNS Benchmark
TABLE IV: Comparison of objective speech evaluation metrics between distilled and undistilled models on the DNS test set.
Methods Casual & NFI Param. (M) FLOPs (G/s) PESQ STOI SISNR (dB)
Noisy - - - 1.58 0.915 9.07
RNNoise [34$\dagger$1 0.1 0.04 1.97 - -
NSNet [35$\dagger$ 1.3 0.13 2.15 0.945 15.61
DTLN [36$\dagger$ 1.0 0.25 - 0.948 16.34
DCCRN [37$\dagger$ 3.7 14.37 2.80 0.964 16.38
FullSubNet+ [38$\dagger$ 8.7 31.12 2.98 0.967 18.34
FRCRN [6$\dagger$ 6.9 12.30 3.23 0.977 19.78
CTS-Net [39$\dagger$ 4.4 5.57 2.94 0.967 17.99
PrimeK-Net [40$\ddagger$ 1.4 44.64 3.34 0.978 19.23
DPDCRN-T 3.5 13.71 3.16 0.973 17.47
DPDCRN-S 0.6 2.44 2.81 0.962 17.05
KD methods
-MSE [32$\ddagger$ 0.6 2.44 2.82 0.962 16.60
-CLSKD [26$\dagger$ 0.6 2.44 2.84 0.963 17.08
-UCLFWPKD [16$\dagger$ 0.6 2.44 2.94 0.966 17.16
-ABC-KD [14$\ddagger$ 0.6 2.44 2.90 0.964 17.14
-Two-Step KD [15$\ddagger$ 0.6 2.44 2.93 0.965 17.05
-M1 (FitNet [33$\ddagger$) 0.6 2.44 2.83 0.962 17.22
-M2 (SemCKD [30$\ddagger$) 0.6 2.44 2.89 0.965 17.36
-M3 (TFCKD) 0.6 2.44 2.97 0.967 17.58
-M4 (I2S-TFCKD) 0.6 2.44 3.03 0.968 17.74
  • 1
    The model results labeled with symbol $\dagger$ are taken directly from the original paper, while the symbol $\ddagger$ denotes the results reproduced on the datasets of this paper according to the original paper’s settings.
Fair comparisons of various algorithms are conducted on the DNS non-reverb test set. Table IV presents the objective speech quality evaluation metrics of various SE methods. We provide details on the causality of each model and whether it uses future frame information (Casual and No Future Information, Casual & NFI). Unlike the 40 ms look-ahead constraint of the DNS challenge, we impose strict restrictions on both the causal architecture and the exclusion of future information to better adapt to real-time applications. Additionally, the model parameter (Param.) count and floating-point operations (FLOPs) per second are provided to evaluate deployment feasibility on edge devices. Among the SE models in Table IV, NSNet, RNNoise, and DTLN are low-complexity real-time SE algorithms primarily based on LSTM networks. These models are fully causal and do not utilize future information. In contrast, DCCRN, FullSubNet+, FRCRN, CTS-Net, and PrimeK-Net fail to satisfy both causality and NFI constraints. Although some architectures, such as DCCRN and CTS-Net, employ causal designs, they still leverage partial future frame information during training and inference. Our distillation baseline DPDCRN, composed of causal convolutions and GRU layers, is strictly causal and operates without any future frame information.
V-B1 Analysis of Distillation Frameworks
Figure 6: The scatter-bubble distribution of PESQ, FLOPs, and model parameters for SE models on the DNS test set.
We compare the effects of different distillation frameworks applied to the base model DPDCRN in Table IV. Among them, MSE that directly narrows the output distance, only achieves slightly higher results than the student model on PESQ and STOI metrics and even falls below the student on the SISNR metric. In contrast, M1 achieves effective improvements compared to output-level distillation through layer-wise feature distillation. On this basis, both CLSKD and M2 incorporate multi-layer knowledge integration to enable cross-layer interaction between the teacher and the student, thus achieving more considerable gains in PESQ and STOI. The distillation advantages of the above multi-layer knowledge transfer frameworks demonstrate that the intermediate representations of the model contain rich transferable information. Recent studies such as UCLFWPKD, ABC-KD, and Two-Step KD design a targeted cross-layer distillation framework according to the characteristics of the SE task, achieving further improvements. In this paper, M3 introduces time-frequency cross-calibration based on the multi-layer distillation framework of M2, achieving notable improvements across all three metrics. This highlights that computing multi-layer calibration weights along both the temporal and spectral domains enables more effective knowledge allocation. Extending M3 to the intra-inter set distillation framework (M4, I2S-TFCKD) achieves the best performance among all distillation methods, demonstrating that knowledge circulation within and between correlated sets further facilitate teacher-student knowledge transfer.
V-B2 Comparison with State-of-the-art SE Methods
Table IV provides the metric performances of current SOTA SE algorithms on the DNS test set. Additionally, Fig. 6 shows the scatter-bubble distribution of PESQ with model parameters and FLOPs. It can be observed that current real-time SE algorithms such as RNNoise, NSNet, and DTLN, despite having extremely low parameters and FLOPs, still exhibit a large performance gap compared to high-complexity SE models. Models like FRCRN and FullSubNet+ have metric advantages, but their high parameter counts (>5M) and FLOPs (>10G) restrict edge-side applicability. The recently proposed SE model PrimeK-Net achieves substantial parameter reduction through targeted architectural design, but its 44.64G FLOPs still impose enormous deployment pressure. This paper applies intra-inter set distillation with time-frequency calibration (I2S-TFCKD) to the student model DPDCRN-S, attaining competitive results across three metrics with low parameter count (0.6M) and computational complexity (2.4G FLOPs).
V-C Comparative Experiments on the L3DAS23 Benchmark
Figure 7: Boxplot of PESQ and the T1 Metric for different distillation strategies on the L3DAS23 validation and development sets.
V-C1 Boxplot Analysis of Distillation Strategies
Fig. 7 presents a boxplot comparison of representative metric results (PESQ and T1 Metric) for different distillation frameworks on the L3DAS23 validation and development sets. From the perspective of the mean value of the metrics, MSE still remains inferior to feature-based distillation strategies, indicating that the utilization of intermediate representation information is also crucial for multi-channel SE distillation. Among all distillation algorithms, the proposed I2S-TFCKD method achieves the highest mean values across both metrics. Compared with the second-highest performing UCLFWPKD, the proposed framework has advantages in both the average metric and the concentration degree of high values, further validating the positive contributions of intra-inter set knowledge flow and time-frequency cross-calibration to the overall distillation effect.
V-C2 Metric comparison Across Models
Table V presents objective metric results on the L3DAS23 development set, comparing multi-channel SE models with and without KD. The base model DPDCRN achieved first place on the blind test set of the L3DAS23 challenge and is causally adjusted in this paper. The teacher backbone DPDCRN-T demonstrates substantial metric advantages over the causal framework EaBNet across all metrics, and ranks second only to the recently proposed non-causal Spatial Net, but the latter still maintains extremely high FLOPs. The compressed student DPDCRN-S exhibits performance degradation compared to the teacher, particularly in PESQ and the WER indicator. Application of various distillation strategies to the student results in varying degrees of improvement across metrics. The proposed I2S-TFCKD method achieves optimal distillation effect, with indicator gains of 0.218 in PESQ, 0.035 in WER, 0.02 in STOI, and 0.028 in the T1 Metric compared to the undistilled student. Overall, the distilled student model maintains causal inference with low computational overhead while achieving competitive performance compared to SOTA SE methods in enhancement results.
TABLE V: Comparison of objective speech evaluation metrics between distilled and undistilled models on the development set in the SE track of L3DAS23.
[…]
Noisy1 - - - 1.166 0.474 0.575 0.550
Neural Beamforming [41$\dagger$ 5.5 32.15 - 0.569 0.684 0.557
EaBNet [42$\dagger$ 2.8 7.38 - 0.549 0.724 0.587
CCA Speech [43$\dagger$ - - - 0.293 0.836 0.771
DeFT-AN [44$\ddagger$ 2.7 37.99 1.976 0.137 0.866 0.864
Spatial Net [45$\ddagger$ 1.6 46.30 2.264 0.133 0.910 0.889
DPDCRN-T 3.5 13.71 2.036 0.129 0.885 0.878
DPDCRN-S 0.6 2.44 1.711 0.185 0.850 0.832
KD methods
-MSE [32$\ddagger$ 0.6 2.44 1.759 0.181 0.847 0.833
-CLSKD [26$\dagger$ 0.6 2.44 1.768 0.178 0.851 0.837
-UCLFWPKD [16$\dagger$ 0.6 2.44 1.852 0.162 0.861 0.850
-ABC-KD [14$\ddagger$ 0.6 2.44 1.795 0.169 0.856 0.843
-Two-Step KD [15$\ddagger$ 0.6 2.44 1.832 0.169 0.859 0.845
-M1 (FitNet [33$\ddagger$) 0.6 2.44 1.774 0.179 0.850 0.836
-M2 (SemCKD [30$\ddagger$) 0.6 2.44 1.851 0.161 0.859 0.849
-M3 (TFCKD) 0.6 2.44 1.914 0.152 0.866 0.857
-M4 (I2S-TFCKD) 0.6 2.44 1.929 0.150 0.870 0.860
  • 1
    Note: Noisy data is taken from the first channel of microphone A.
VI Conclusion
In this paper, we proposed an intra-inter set distillation framework with time-frequency calibration (I2S-TFCKD) for SE model compression. Specifically, I2S-TFCKD partitions intermediate layers of SE models into multiple correlated sets. Within each set, intra-set cross-layer knowledge transfer is performed, and representative features are generated through residual fusion. The representative features of multiple correlated sets are then integrated into a fused feature set to enable inter-set distillation. Furthermore, we designed a multi-layer time-frequency cross-calibration distillation tailored to SE tasks: the temporal stream captures importance distributions across speech frames, while the spectral stream focuses on distillation contributions of layer-wise spectral features. We verified the effectiveness of time-frequency calibration and intra-inter set distillation through ablation studies. Experimental results on two public datasets (DNS and L3DAS23) demonstrated that the proposed distillation framework stably and effectively improves the enhancement performance of the low-complexity student model, narrows the gap with the teacher and outperforms other distillation strategies across various metrics. Future research will focus on multi-teacher distillation for SE models, enabling an elite student to handle various acoustic scenarios through knowledge integration from multiple teachers.

### v2

@@
-I2S-TFCKD: Intra-Inter Set Knowledge Distillation with Time-Frequency Calibration for Speech Enhancement
+
 Abstract
 In recent years, complexity compression of neural network (NN)-based speech enhancement (SE) models has gradually attracted the attention of researchers, especially in scenarios with limited hardware resources or strict latency requirements.
-The main difficulties and challenges lie in achieving a balance between complexity and performance according to the characteristics of the task.
-In this paper, we propose an intra-inter set knowledge distillation (KD) framework with time-frequency calibration (I2S-TFCKD) for SE.
+In this paper, we propose an intra-inter representation fusion knowledge distillation (KD) framework with time-frequency calibration (I2RF-TFCKD) for SE, which achieves distillation through the fusion of multi-layer teacher-student feature flows.
 Different from previous distillation strategies for SE, the proposed framework fully utilizes the time-frequency differential information of speech while promoting global knowledge flow.
-Firstly, we propose a multi-layer interactive distillation based on dual-stream time-frequency cross-calibration, which calculates the teacher-student similarity calibration weights in the time and frequency domains respectively and performs cross-weighting, thus enabling refined allocation of distillation contributions across different layers according to speech characteristics.
-Secondly, we construct a collaborative distillation paradigm for intra-set and inter-set correlations.
+Firstly, we construct a collaborative distillation paradigm for intra-set and inter-set correlations.
 Within a correlated set, multi-layer teacher-student features are pairwise matched for calibrated distillation.
 Subsequently, we generate representative features from each correlated set through residual fusion to form the fused feature set that enables inter-set knowledge interaction.
+Secondly, we propose a multi-layer interactive distillation based on dual-stream time-frequency cross-calibration, which calculates the teacher-student similarity calibration weights in the time and frequency domains respectively and performs cross-weighting, thus enabling refined allocation of distillation contributions across different layers according to speech characteristics.
 The proposed distillation strategy is applied to the dual-path dilated convolutional recurrent network (DPDCRN) that ranked first in the SE track of the L3DAS23 challenge.
-To evaluate the effectiveness of I2S-TFCKD, we conduct experiments on both single-channel and multi-channel SE datasets.
+To evaluate the effectiveness of I2RF-TFCKD, we conduct experiments on both single-channel and multi-channel SE datasets.
 Objective evaluations demonstrate that the proposed KD strategy consistently and effectively improves the performance of the low-complexity student model and outperforms other distillation schemes.
-Index Terms:
-speech enhancement, knowledge distillation, time-frequency calibration, intra-inter set correlations, residual fusion.
-publicationid: pubid: $0$–$00$ 2025 IEEE
-I Introduction
+The codes are available at https://github.com/JMCheng-SEU/I2RF-TFCKD-SE.
+keywords:
+[…]
+1 Introduction
 Speech enhancement (SE) aims to remove background noises and recover clean speech from noisy mixtures.
@@
 [40].
-Traditional SE methods based on statistical signal processing [2] estimate the noise spectrum from noisy speech for suppression, requiring low computational resources.
-However, these algorithms typically rely on stationarity assumptions, which fundamentally constrain their performance in handling non-stationary noises encountered in real-world scenarios [3].
 With the recent emergence of data-driven deep learning approaches, SE models have begun leveraging massive training data to characterize both speech signals, noise components, and their nonlinear relationships.
@@
 Consequently, the exploration of SE models balancing performance with lightweight design has emerged as a research hotspot and trend.
-Model compression is an effective solution, including pruning [7], quantization [8], and knowledge distillation (KD) [9].
-Stamenovic et al.
-[10] specifically investigate pruning methodologies tailored for SE frameworks.
+Model compression is an effective solution, including pruning, quantization, and knowledge distillation (KD).
 Tan et al.
@@
 Kim et al.
-[12] uses the teacher’s pseudo target to train the student model, enabling a zero-shot learning procedure.
+[13] use the teacher’s pseudo target to train the student model, enabling a zero-shot learning procedure.
 Thakker et al.
@@
 Knowledge transfer of intermediate model representations has gradually attracted researchers’ attention, with the key challenge lying in the SE-specific distillation design.
-In our previous work [16], we proposed a unified residual fusion probabilistic KD framework for SE, which leverages cross-layer supervision to guide information transfer between teacher and student while introducing a frame weighting probabilistic distillation loss to perceive key frame information.
-In this paper, we design an intra-inter set KD with time-frequency calibration (I2S-TFCKD) framework to further optimize the cross-layer supervision distillation paradigm.
-On the one hand, we introduce temporal and spectral domain cross-computation to derive multi-layer distillation calibration weights, leveraging time-frequency characteristics for targeted distillation optimization.
-On the other hand, unlike previous approaches that conduct cross-layer interactions within isolated correlation sets, we propose collaborative distillation across intra-set and inter-set correlations, effectively promoting knowledge flow throughout the model architecture.
+In this paper, we design an intra-inter representation fusion with time-frequency calibration KD (I2RF-TFCKD) framework to further optimize the cross-layer supervision distillation paradigm.
+On the one hand, unlike previous approaches [21, 4] that conduct cross-layer interactions within isolated correlation sets, we propose collaborative distillation across intra-set and inter-set correlations, effectively promoting knowledge flow throughout the model architecture.
+On the other hand, we introduce temporal and spectral domain cross-computation to derive multi-layer distillation calibration weights, leveraging time-frequency characteristics for targeted distillation optimization.
 The backbone network for distillation employs our previously proposed dual-path dilated convolutional recurrent network (DPDCRN) model [5], which won the SE track of the L3DAS23 challenge.
 The main contributions of this work are as follows:
-  • We propose multi-layer time-frequency cross-calibration distillation, which independently computes teacher-student similarity calibration weights along temporal and spectral dimensions for cross-weighting, enabling refined allocation of distillation contributions across different layers.
   • We design a collaborative distillation paradigm for intra-set and inter-set correlations.
-First, we partition the SE model structure into multiple correlation sets for intra-set multi-layer teacher-student matching.
-Subsequently, representative features from each correlated set are generated through residual recursive fusion, forming the fused feature set that enables inter-set knowledge interaction.
-  • Extensive evaluations on public SE benchmarks demonstrate that our proposed I2S-TFCKD framework for SE surpasses existing distillation methods and enabling the student model to achieve competitive enhancement results with low computational complexity.
-II Related Work
-In this section, we provide an overview of related works on distillation methodologies for SE, particularly focusing on feature distillation-related research.
-II-A Knowledge Distillation Strategy
-KD originated from the pioneering work of Hinton et al.
-[18], where the student model learns from both ground-truth labels and soft labels provided by the teacher model.
-Building upon this, subsequent research focuses on exploring effective regularization methods to balance soft and hard labels.
-In recent years, feature distillation has gradually come into the spotlight, with numerous studies demonstrating that distilling intermediate feature representations exhibits stronger generalization ability and more promising results compared to soft-label distillation.
-Niu et al.
-[19] transform the feature maps of different levels into a frequency-enhanced feature space, which highlights the low-frequency features for both the teacher and student networks.
-Cross-image relational knowledge extraction is proposed in [20], leveraging pixel-to-pixel and pixel-to-region distillation to fully utilize the structural relationships between images.
-These methods primarily utilize single-layer features for knowledge transfer, while recent advances explore multi-layer and multi-stage feature distillation frameworks.
-A cross-layer knowledge review mechanism is investigated in [21], which integrates multi-layer teacher features to guide the student.
-Zhao et al.
-[22] developed a dense KD in all aspects manner, using the dynamic class-specific pseudo-labeling strategy to distill intermediate features and output logits collaboratively with more emphasis on semantic-invariant knowledge inheritance.
-Overall, feature distillation can transfer fine-grained knowledge such as edge, texture, and semantic features, garnering increasing research attention.
-Cross-layer feature distillation transcends traditional layer-wise correspondence constraints, allowing a student to simultaneously learn low-level detail features and high-level abstract representations from teachers.
-However, this paradigm also introduces challenges in representation alignment, making it a worthwhile direction for in-depth research.
-II-B Distillation for Speech Enhancement
-Most existing distillation works discussed above have been primarily applied to the computer vision and natural language processing domains.
-In the acoustic field, KD methods are initially adopted for classification tasks such as acoustic scene classification [23] and automatic speech recognition [24].
-Subsequently, KD for regression tasks like SE has gradually emerged.
+Specifically, the SE model is partitioned into multiple correlated sets, and multi-layer teacher–student matching distillation is performed both within each set and across sets.
+  • We introduce a recursive fusion mechanism to generate representative features for each correlated set from both teacher and student, and then integrate them into a fused feature set.
+The recursive fusion preserves the critical information of each set while reducing redundancy in cross-set pairing.
+  • We propose multi-layer time-frequency cross-calibration, which independently computes teacher-student similarity calibration weights along temporal and spectral dimensions for cross-weighting, enabling refined allocation of distillation contributions across different layers.
+  • Extensive evaluations on public SE benchmarks demonstrate that our proposed I2RF-TFCKD framework for SE surpasses existing distillation methods and enabling the student model to achieve competitive enhancement results with low computational complexity.
+The remainder of the paper is organized as follows: Section 2 introduces related works.
+Section 3 presents a detailed description of the proposed framework.
+The experimental setup is outlined in Section 4, while Section 5 covers the experimental results and discussion.
+Finally, conclusions are presented in Section 6.
+2 Related work
+2.1 Time-frequency processing for SE
+Speech signals inherently exhibit temporal evolution characteristics (e.g., syllabic rhythm and prosody) and frequency properties (e.g., formant and harmonic structures).
+Time-domain processing captures the nonuniform variations of speech energy along the temporal axis, whereas frequency-domain representations are well suited to differentiating the distribution patterns of speech and noise.
+Consequently, the joint processing of time and frequency domains (TF processing) has received increasing attention in recent SE research [40].
+On one hand, TF processing has been deeply integrated into the architectural design of SE networks.
+Models based on convolutional–recurrent frameworks [12, 39] utilize convolutional modules to extract local spectral patterns in the frequency domain, while recurrent units are used to capture temporal dependencies in the time domain.
+More recent designs [14, 22] adopt parallel time–frequency dual-path branches, unfolding intermediate representations sequentially along the time and frequency axes and employing recurrent units to alternately process and integrate information across both domains.
+All these architectures explicitly emphasize the interaction between time and frequency domains, jointly preserving temporal continuity and spectral details.
+On the other hand, TF interaction has also been reflected in the design of loss functions for SE models.
+Xia et al.
+[36] introduce two mean-squared-error (MSE)-based spectral losses, where frame-level voice activity detection (VAD) in the time domain is used to separate the speech distortion and noise suppression terms.
+The modulation-domain loss [30], formulated via the spectro-temporal modulation index (STMI), evaluates the integrity of processed speech within a perceptually motivated TF modulation space.
+These TF domain losses fully leverage speech-specific characteristics to formulate regression distance measures.
+Overall, TF processing enables SE models to simultaneously capture temporal context and perceive fine spectral structures, thereby providing a promising direction for optimizing SE frameworks.
+2.2 Knowledge distillation for SE
+Knowledge distillation [11] has been widely applied in the fields of computer vision and natural language processing.
+In the acoustic field, KD methods are initially adopted for classification tasks such as automatic speech recognition [37].
+Subsequently, KD for regression tasks like text-to-speech (TTS) and SE has gradually emerged [19].
 Early distillation studies for SE models primarily focus on minimizing output-level discrepancies between teachers and students, exemplified by elite sub-band distillation [10].
@@
 A common challenge in representation-based distillation arises from dimension mismatch between teacher and student features.
-To address this, our previous work [26] first proposes a cross-layer similarity distillation framework for SE models, which aligns teacher-student features via similarity loss.
-Similarly, multiple studies have adopted multi-layer distillation strategies to compress SE models, including multi-view attention transfer distillation [27] and two-step fine-grained similarity-preserving KD [15].
+To address this, a cross-layer similarity distillation framework [3] is proposed for SE models, which aligns teacher-student features via similarity loss.
+Similarly, multiple studies have adopted multi-layer distillation strategies to compress SE models, including bidirectional KD [1] and two-step fine-grained similarity-preserving KD [21].
 It is worth noting that recent studies have begun to focus on leveraging the inherent characteristics of the SE task to improve distillation frameworks.
@@
 To address these limitations, this work mainly includes two innovative improvements: the time-frequency calibration strategy for distillation and the intra-inter set global knowledge transfer mechanism.
-III Methodology
-III-A Teacher and Student Network Architecture
+3 Methodology
+3.1 Teacher and student network architecture
 Figure 1: Backbone network architecture of the teacher model.
-Figure 2: Overall architecture of the I2S-TFCKD framework.
+In this paper, we extend the multi-layer distillation within correlated sets [21, 4] to intra-inter set global knowledge transfer, while implementing time-frequency cross-assignment weighting in the distillation loss.
+The backbone architecture chooses DPDCRN [5] that won the SE track of the L3DAS23 challenge, due to its facility for structural compression and scalability to both single-channel and multi-channel SE tasks.
+The backbone architecture of the teacher model is presented in Fig.
+1, where $B$ is the batch size, $C$ is the number of output channels, $T$ is the number of speech frames, $D$ is the feature dimension size.
+The input of the DPDCRN model is the complex spectrum obtained by applying the short-time Fourier transform (STFT) to the noisy speech.
+The output is the estimated complex ratio mask, which is multiplied with the noisy spectrum in the complex domain, followed by an inverse STFT (iSTFT) to reconstruct the enhanced speech.
+The main structure of DPDCRN consists of a convolutional encoder, a convolutional decoder, and an intermediate frequency-time (F-T) processing module.
+The encoder contains two convolutional layers and four dilated convolutional blocks, while the decoder contains two deconvolutional layers and four dilated convolutional blocks.
+Each frequency-time processing module contains dual branches for frequency and temporal axes, with each branch comprising a causal multi-head self-attention (MHSA) module and a feedforward network based on gated recurrent units (GRUs).
+Layer normalization (LNorm) and ReLU activation are incorporated between layers to regulate the distribution of feature flows.
+For each convolutional layer, the student has half the number of channels of the teacher.
+For the F-T processing modules, the number of hidden units in the student is half that of the teacher.
+Additionally, the teacher model has four cascaded F-T modules, while the student only uses one.
+Overall, the student contains merely 17% of the teacher’s parameters.
+Figure 2: Overall architecture of the I2RF-TFCKD framework.
 We present the detailed process of intra-inter set distillation.
@@
 2(c) visualizes the process of generating fused features by utilizing the features of the current layer and the inherited recursive features.
-In our previous work [16], we designed a residual fusion distillation framework for SE based on the DPDCRN architecture that won the SE track of L3DAS23 challenge.
-In this paper, we extend the multi-layer distillation within correlated sets to intra-inter set global knowledge transfer, while implementing time-frequency cross-assignment weighting in the distillation loss.
-The backbone architecture remains DPDCRN due to its facility for structural compression and scalability to both single-channel and multi-channel SE tasks.
-The backbone architecture of the teacher model is presented in Fig.
-1.
-The main structure of DPDCRN consists of a convolutional encoder, a convolutional decoder, and an intermediate frequency-time (F-T) processing module.
-The encoder contains two convolutional layers and four dilated convolutional blocks, while the decoder contains two deconvolutional layers and four dilated convolutional blocks.
-Each frequency-time processing module contains dual branches for frequency and temporal axes, with each branch comprising a causal multi-head self-attention (MHSA) module and a feedforward network based on gated recurrent units (GRUs).
-For each convolutional layer, the student has half the number of channels of the teacher.
-For the F-T processing modules, the number of hidden units in the student is half that of the teacher.
-Additionally, the teacher model has four cascaded F-T modules, while the student only uses one.
-Overall, the student contains merely 17% of the teacher’s parameters.
-III-B Intra-inter Set Knowledge Transfer Architecture
+3.2 Intra-inter representation fusion knowledge transfer architecture
 According to the encoder-decoder framework of current mainstream SE models, we partition all layers of teacher-student networks into three correlated sets corresponding to the encoder, decoder, and intermediate F-T processing modules, respectively, to achieve hierarchical knowledge transfer.
-In previous work [16], we utilized residual fusion to implement cross-layer knowledge transfer within each correlated set and confirmed its advantages over conventional layer-wise distillation.
-However, residual fusion distillation introduces redundant replication operations when the number of modules in teacher and student networks differs [16], and cross-set information interaction remains unexplored.
+In previous work [4], residual fusion is utilized to implement cross-layer knowledge transfer within each correlated set and confirmed its advantages over conventional layer-wise distillation.
+However, residual fusion distillation introduces redundant replication operations when the number of modules in teacher and student networks differs, and cross-set information interaction remains unexplored.
 In this paper, we propose a unified cross-layer distillation framework that jointly addresses intra-set and inter-set knowledge transfer.
@@
 2.
-III-C Time-frequency Cross-calibrated Multi-layer Matching Distillation
-The candidate correlation set of the teacher and student is denoted as ${\cal P}=\left\{{\left({{l_{s}},{l_{t}}}\right)\left|{\forall{l_{s}}\in\left[{1,\ldots,{L_{s}}}\right]}\right.,{l_{t}}\in\left[{1,\ldots,{L_{t}}}\right]}\right\}$, where ${{l_{s}}}$ and ${{l_{t}}}$ represent the corresponding layers of the candidate student and teacher, respectively.
-In the feature distillation paradigm, the feature maps of candidate teacher-student pairs are first transformed into specific embedding representations, and then, the distance between teacher-student embeddings is calculated as the distillation loss ${{\cal L}_{KD}}$:
-${{\cal L}_{KD}}=\sum\limits_{\left({{l_{s}},{l_{t}}}\right)\in{\cal P}}{{\cal D}\left({Tran{s^{s}}\left({{{\bf{F}}_{{l_{s}}}}}\right),Tran{s^{t}}\left({{{\bf{F}}_{{l_{t}}}}}\right)}\right)},$ (1)
+3.2.1 Intra-set multi-layer teacher–student feature distillation
+The candidate teacher-student correlation set is denoted as ${\cal P}=\left\{{\left({{l_{s}},{l_{t}}}\right)\left|{\forall{l_{s}}\in\left[{1,\ldots,{L_{s}}}\right]}\right.,{l_{t}}\in\left[{1,\ldots,{L_{t}}}\right]}\right\}$, where $l_{s}$}} and $l_{t}$}} represent the corresponding layers of the candidate student and teacher, respectively.
+In the feature distillation paradigm, the feature maps of candidate teacher-student pairs are first transformed into specific embedding representations, and then, the distance between teacher-student embeddings is calculated as the distillation loss.
+As shown in Fig.
+3(a), single-layer feature distillation only transfers knowledge between teacher and student at the same hierarchical level:
+${{\cal L}_{singleKD}}=\sum\limits_{i\in{\cal P}}{{\cal D}\left({Tran{s^{s}}\left({{{\bf{F}}_{l_{s}^{i}}}}\right),Tran{s^{t}}\left({{{\bf{F}}_{l_{t}^{i}}}}\right)}\right)},$ (1)
 where ${\bf{F}}_{{l_{s}}}$} and ${\bf{F}}_{{l_{t}}}$} denote the hidden layer feature representations of the student and teacher respectively, $Tran{s^{s}}$ and $Tran{s^{t}}$ are used to transform teacher-student feature pairs into specific embedding representations for alignment.
-Conventional layer-wise distillation imposes strict constraints on structural and layer alignment between teacher and student models, requiring redundant layer information to be discarded in cases of mismatch.
-In contrast, multi-layer distillation that matches single-layer student features with multi-layer teacher features, is not constrained by the difference in the number of layers.
-Different layer structures in DNNs exhibit varying capabilities to learn feature representations.
-As the depth of layers increases, the intermediate representation information of the model gradually becomes more abstract.
-Although multi-layer feature matching enables student models to perceive a broader range of teacher hidden knowledge, indiscriminate alignment may cause semantic redundancy that degrades learning efficiency and performance.
-To maximize absorption of valuable information, student layers should preferentially align with teacher layers exhibiting high semantic relevance.
-Inspired by the observation that the proximity of pairwise similarity matrices can be regarded as a good measurement of the inherent semantic similarity [30], we use semantic-aware similarity for distillation calibration.
-Specifically, we introduce a calibration weight ${\alpha_{\left({{l_{s}},{l_{t}}}\right)}}$ to dynamically adjust the contribution of candidate teacher-student layer pairs based on their semantic congruence, and constrain the corresponding weights through a softmax function: $\sum\limits_{{l_{t}}=1}^{{L_{t}}}{{\alpha_{\left({{l_{s}},{l_{t}}}\right)}}}=1,\forall{l_{s}}\in\left[{1,\ldots,{L_{s}}}\right]$.
-For SE models, time-frequency interaction analysis enables simultaneous capture of local spectral variations and temporal dynamics in speech signals, which is expected to facilitate more refined alignment of student and correlated teachers.
-Therefore, we compute calibration weights for temporal and spectral flows independently and perform cross-weighting.
-Given a batch of feature map input ${{\bf{F}}_{s/t}}\in{{\mathop{\rm R}\nolimits}^{B\times C\times T\times D}}$, where $B$ is the batch size, $C$ is the number of output channels, $T$ is the number of speech frames, $D$ is the feature dimension size, and $s/t$ denotes the student or teacher, we first calculate the similarity maps in the time-flow and frequency-flow respectively, as shown in Fig.
-3.
-For the time-domain flow, the feature map is split into multiple instances along the batch dimension, then flattened along the channel and feature dimensions.
-${l^{2}}$ Normalization is applied to the merged feature dimension to obtain the temporal transformed features ${\bf{M}}_{s/t}^{T}\in{{\mathop{\rm R}\nolimits}^{B\times T\times\left({C\times D}\right)}}$.
-Finally, the time-flow mapping matrix ${\bf{P}}_{s/t}^{T}\in{{\mathop{\rm R}\nolimits}^{B\times T\times T}}$ is computed using cosine similarity measurement:
-${\bf{P}}_{s/t}^{T}=\frac{1}{2}\left({\frac{{{{\left({{\bf{M}}_{s/t}^{T}}\right)}^{\mathop{\rm T}\nolimits}}\cdot{\bf{M}}_{s/t}^{T}}}{{\left\|{{\bf{M}}_{s/t}^{T}}\right\|_{2}^{2}}}+1}\right).$ (2)
-For the frequency-domain flow, we partition the feature maps along the time dimension into multiple instances to maintain frame independence.
-Then, the frequency transformed features ${\bf{M}}_{s/t}^{F}\in{{\mathop{\rm R}\nolimits}^{T\times B\times\left({C\times D}\right)}}$ are obtained by flattening and normalization.
-Finally, the frequency-flow mapping matrix ${\bf{P}}_{s/t}^{F}\in{{\mathop{\rm R}\nolimits}^{T\times B\times B}}$ is also calculated by cosine similarity.
-Subsequently, we employ trainable weight matrices to project the teacher-student self-similarity mapping matrices from both the temporal and spectral domains into query and key subspaces, mitigating the effects of noise and sparsity:
-${\bf{Q}}_{{l_{s}}}^{T/F}=EM{B_{Q}}\left({{\bf{P}}_{{l_{s}}}^{T/F}}\right){\rm{,}}$ (3)
-${\bf{K}}_{{l_{t}}}^{T/F}=EM{B_{K}}\left({{\bf{P}}_{{l_{t}}}^{T/F}}\right),$
-where $EM{B_{Q}}\left(\cdot\right)$ and $EM{B_{K}}\left(\cdot\right)$ consist of the fully connected layer, ReLU activation, and layer normalization.
-The time-domain and frequency-domain flows adopt independent trainable weight parameters.
-The similarity calibration coefficients for the time and frequency domains are computed through domain-specific query-key interactions:
-$\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{T/F}={e^{{{\left({{\bf{Q}}_{{l_{s}}}^{T/F}}\right)}^{\rm{T}}}{\bf{K}}_{{l_{t}}}^{T/F}}}\cdot{\left({\sum\limits_{{l_{t}}\in{\cal P}}{{e^{{{\left({{\bf{Q}}_{{l_{s}}}^{T/F}}\right)}^{\rm{T}}}{\bf{K}}_{{l_{t}}}^{T/F}}}}}\right)^{-1}}.$ (4)
-Finally, cross-layer matching is performed for multiple teacher-student feature pairs within the current correlated set, where similarity metrics are computed through both time and frequency flows and distillation contributions are allocated via $T/F$ calibration coefficients.
-The multi-layer time-frequency calibration knowledge distillation (TFCKD) loss is calculated as follows:
-${\cal L}_{{\rm{TFCKD}}}$ $=\sum\limits_{{l_{s}}=1}^{{L_{s}}}\sum\limits_{{l_{t}}=1}^{{L_{t}}}\left(\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{T}\left({{\bf{P}}_{{l_{t}}}^{T}-{\bf{P}}_{{l_{s}}}^{T}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{T}}}{{{\bf{P}}_{{l_{s}}}^{T}}}}\right)\right.$ (5)
-$\left.+\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{F}\left({{\bf{P}}_{{l_{t}}}^{F}-{\bf{P}}_{{l_{s}}}^{F}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{F}}}{{{\bf{P}}_{{l_{s}}}^{F}}}}\right)\right),$
-where ${L_{s}}$ and ${L_{t}}$ are the total number of layers of the student and teacher in the current correlated set.
-This distillation loss serves as the knowledge transfer mechanism for both intra-set and inter-set teacher-student features.
-Figure 3: Similarity mapping of time and frequency flows.
-The self-similarity matrices for distillation are calculated in two flows: the time domain and the frequency domain.
-The distribution of features along the time axis will affect the similarity computation in the time flow, whereas the frequency domain processing operates frame-wise.
-III-D Residual Fusion Features for Correlated Sets
+$\cal D$} is the computation of distillation distance.
+However, layer-wise distillation imposes strict constraints on structural and layer alignment between teacher and student models, requiring redundant layer information to be discarded in cases of mismatch.
+In addition, one-to-one knowledge transfer makes it difficult for the student model to progressively learn hierarchical information from the teacher.
+Therefore, we introduce multi-layer matching distillation for knowledge transfer within correlated sets.
+As illustrated in Fig.
+3(b), The single-layer student feature is aligned with hierarchical teacher features within the correlated set:
+${{\cal L}_{multiKD}}=\sum\limits_{\left({{l_{s}},{l_{t}}}\right)\in{\cal P}}{{\cal D}\left({Tran{s^{s}}\left({{{\bf{F}}_{{l_{s}}}}}\right),Tran{s^{t}}\left({{{\bf{F}}_{{l_{t}}}}}\right)}\right)}.$ (2)
+Nevertheless, cross-layer distillation may incur redundant computations.
+To address this, we assign distillation weights across layers according to semantic information, which will be further discussed in Section 3.3.
+Figure 3: (a) using single-level teacher knowledge to guide one-level learning of the student. (b) using multiple layers of the teacher to supervise one layer in the student.
+3.2.2 Inter-set distillation based on recursive representation fusion
 We partition the teacher-student correlated sets based on the functionality of structural blocks in the SE model.
@@
 ${{\bf{T}}^{j}}={\rm{2DConv}}\left({{{\bf{R}}^{j}}}\right).$ (6)
-For each correlated set, we select the last recursion output as the representative fusion feature of the current set, which is then integrated to form a fused feature set.
+For each correlated set, we select the last recursion output as the representative fused feature of the current set, which is then integrated to form a fused feature set.
 We believe that due to the symmetry of the encoder-decoder architecture in SE models, higher-level features from the middle layers exhibit stronger capability to learn effective information from lower-level features near the input/output ends.
@@
 Corresponding representative features are generated for each correlated set in both the teacher and student models following the above process, and integrated into the teacher-student fused feature set for inter-set distillation.
-III-E Training Procedure
-In this section, we elaborate on the entire training procedure for intra-inter set distillation.
+3.3 Time-frequency cross calibration for multi-layer hierarchical matching
+Figure 4: Similarity mapping of time and frequency flows.
+The self-similarity matrices for distillation are calculated in two flows: the time domain and the frequency domain.
+The distribution of features along the time axis will affect the similarity computation in the time flow, whereas the frequency domain processing operates frame-wise.
+Different layer structures in DNNs exhibit varying capabilities to learn feature representations.
+As the depth of layers increases, the intermediate representation information of the model gradually becomes more abstract.
+Although multi-layer feature matching enables student models to perceive a broader range of teacher hidden knowledge, indiscriminate alignment may cause semantic redundancy that degrades learning efficiency and performance.
+To maximize absorption of valuable information, student layers should preferentially align with teacher layers exhibiting high semantic relevance.
+Inspired by the observation that the proximity of pairwise similarity matrices can be regarded as a good measurement of the inherent semantic similarity [32], we use semantic-aware similarity for distillation calibration.
+Specifically, we introduce a calibration weight $\alpha_{\left({{l_{s}},{l_{t}}}\right)}$} to dynamically adjust the contribution of candidate teacher-student layer pairs based on their semantic congruence, and constrain the corresponding weights through a softmax function: $\sum\limits_{{l_{t}}=1}^{{L_{t}}}{{\alpha_{\left({{l_{s}},{l_{t}}}\right)}}}=1,\forall{l_{s}}\in\left[{1,\ldots,{L_{s}}}\right]$.
+For SE models, time-frequency interaction analysis enables simultaneous capture of local spectral variations and temporal dynamics in speech signals, which is expected to facilitate more refined alignment of student and correlated teachers.
+Therefore, we compute calibration weights for temporal and spectral flows independently and perform cross-weighting.
+Given a batch of feature map input ${{\bf{F}}_{s/t}}\in{{\mathop{\rm R}\nolimits}^{B\times C\times T\times D}}$, where $B$ is the batch size, $C$ is the number of output channels, $T$ is the number of speech frames, $D$ is the feature dimension size, and $s/t$ denotes the student or teacher, we first calculate the similarity maps in the time-flow and frequency-flow respectively, as shown in Fig.
+4.
+For the time-domain flow, the feature map is split into multiple instances along the batch dimension, then flattened along the channel and feature dimensions.
+$l^{2}$} Normalization is applied to the merged feature dimension to obtain the temporal transformed features ${\bf{M}}_{s/t}^{T}\in{{\mathop{\rm R}\nolimits}^{B\times T\times\left({C\times D}\right)}}$.
+Finally, the time-flow mapping matrix ${\bf{P}}_{s/t}^{T}\in{{\mathop{\rm R}\nolimits}^{B\times T\times T}}$ is computed using cosine similarity measurement:
+${\bf{P}}_{s/t}^{T}=\frac{1}{2}\left({\frac{{{{\left({{\bf{M}}_{s/t}^{T}}\right)}^{\mathop{\rm T}\nolimits}}\cdot{\bf{M}}_{s/t}^{T}}}{{\left\|{{\bf{M}}_{s/t}^{T}}\right\|_{2}^{2}}}+1}\right).$ (7)
+For the frequency-domain flow, we partition the feature maps along the time dimension into multiple instances to maintain frame independence.
+Then, the frequency transformed features ${\bf{M}}_{s/t}^{F}\in{{\mathop{\rm R}\nolimits}^{T\times B\times\left({C\times D}\right)}}$ are obtained by flattening and normalization.
+Finally, the frequency-flow mapping matrix ${\bf{P}}_{s/t}^{F}\in{{\mathop{\rm R}\nolimits}^{T\times B\times B}}$ is also calculated by cosine similarity.
+Subsequently, we employ trainable weight matrices to project the teacher-student self-similarity mapping matrices from both the temporal and spectral domains into query and key subspaces, mitigating the effects of noise and sparsity:
+${\bf{Q}}_{{l_{s}}}^{T/F}=EM{B_{Q}}\left({{\bf{P}}_{{l_{s}}}^{T/F}}\right){\rm{,}}$ (8)
+${\bf{K}}_{{l_{t}}}^{T/F}=EM{B_{K}}\left({{\bf{P}}_{{l_{t}}}^{T/F}}\right),$
+where $EM{B_{Q}}\left(\cdot\right)$ and $EM{B_{K}}\left(\cdot\right)$ consist of the fully connected layer, ReLU activation, and layer normalization.
+The time-domain and frequency-domain flows adopt independent trainable weight parameters.
+The similarity calibration coefficients for the time and frequency domains are computed through domain-specific query-key interactions:
+$\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{T/F}={e^{{{\left({{\bf{Q}}_{{l_{s}}}^{T/F}}\right)}^{\rm{T}}}{\bf{K}}_{{l_{t}}}^{T/F}}}\cdot{\left({\sum\limits_{{l_{t}}\in{\cal P}}{{e^{{{\left({{\bf{Q}}_{{l_{s}}}^{T/F}}\right)}^{\rm{T}}}{\bf{K}}_{{l_{t}}}^{T/F}}}}}\right)^{-1}}.$ (9)
+Finally, cross-layer matching is performed for multiple teacher-student feature pairs within the current correlated set, where similarity metrics are computed through both time and frequency flows and distillation contributions are allocated via $T/F$ calibration coefficients.
+The multi-layer time-frequency calibration knowledge distillation (TFCKD) loss is calculated as follows:
+${\cal L}_{{\rm{TFCKD}}}=\sum\limits_{{l_{s}}=1}^{{L_{s}}}\sum\limits_{{l_{t}}=1}^{{L_{t}}}\left(\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{T}\left({{\bf{P}}_{{l_{t}}}^{T}-{\bf{P}}_{{l_{s}}}^{T}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{T}}}{{{\bf{P}}_{{l_{s}}}^{T}}}}\right)+\alpha_{\left({{l_{s}},{l_{t}}}\right)}^{F}\left({{\bf{P}}_{{l_{t}}}^{F}-{\bf{P}}_{{l_{s}}}^{F}}\right)\log\left({\frac{{{\bf{P}}_{{l_{t}}}^{F}}}{{{\bf{P}}_{{l_{s}}}^{F}}}}\right)\right),$ (10)
+where $L_{s}$} and $L_{t}$} are the total number of layers of the student and teacher in the current correlated set.
+This distillation loss serves as the knowledge transfer mechanism for both intra-set and inter-set teacher-student features.
+3.4 Training procedure
+Input: Student and teacher correlated sets $\left\{{C_{s}^{k}}\right\}_{k=1}^{q}$ and $\left\{{C_{t}^{k}}\right\}_{k=1}^{q}$.
+Output: Intra-set and inter-set distillation loss ${\cal L}_{Intra}$} and ${\cal L}_{Inter}$}.
+1 for $k\leftarrow 1$ to $q$ do /* Iterate through all correlated sets. */
+2  Sample intra-set student representations $\left({{\bf{F}}_{s}^{\rm{1}},{\bf{F}}_{s}^{2}\ldots,{\bf{F}}_{s}^{m}}\right)$ from $C_{s}^{k}$;
+3  Sample intra-set teacher representations $\left({{\bf{F}}_{t}^{\rm{1}},{\bf{F}}_{t}^{2}\ldots,{\bf{F}}_{t}^{n}}\right)$ from $C_{t}^{k}$;
+4  ${{\cal L}_{Intra}}=0$;
+5
+6 for $i\leftarrow 1$ to $m$ do /* Iterate through all teacher-student layers in the current set. */
+7     for $j\leftarrow 1$ to $n$ do
+8        ${{\cal L}_{Intra}}+={\rm{TF\_Calibration}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)$;
+9        if $i==m$ then /* Ensure the uniqueness. */
+10           if $j==1$ then
+11              ${\bf{R}}_{t}^{j}={\bf{F}}_{t}^{j}$;
+12
+13          else
+14              ${\bf{T}}_{t}^{j},{\bf{R}}_{t}^{j}={\rm{Fusion}}\left({{\bf{F}}_{t}^{j},{\bf{R}}_{t}^{j-1}}\right)$;
+15
+16           end if
+17
+18        end if
+19
+20     end for
+21    if $i==1$ then
+22        ${\bf{R}}_{s}^{j}={\bf{F}}_{s}^{j}$;
+23
+24    else
+25        ${\bf{T}}_{s}^{j},{\bf{R}}_{s}^{j}={\rm{Fusion}}\left({{\bf{F}}_{s}^{j},{\bf{R}}_{s}^{j-1}}\right)$;
+26
+27     end if
+28
+29  end for
+ /* The final output is chosen as the representative fused feature. */
+30  ${\bf{u}}_{s}^{k}={\bf{T}}_{s}^{m}$;
+31  ${\bf{u}}_{t}^{k}={\bf{T}}_{t}^{n}$;
+32
+33 end for
+[…]
+35
+36for $i\leftarrow 1$ to $q$ do /* Iterate through fused features. */
+37  for $j\leftarrow 1$ to $q$ do
+38     ${{\cal L}_{Inter}}+={\rm{TF\_Calibration}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)$;
+39
+40  end for
+41
+42 end for
+Algorithm 1 Intra-inter distillation processing
+In this section, we elaborate on the entire training procedure for intra-inter distillation.
 The overall workflow is outlined in Algorithm 1.
@@
 The total loss for training the student model is formulated as:
-${{\cal L}_{stu}}$ $={{\cal L}_{MRSTFT}}+\sum\limits_{k=1}^{q}{\sum\limits_{i=1}^{m}{\sum\limits_{j=1}^{n}{{{\cal L}_{Intra}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)}}}$ (10)
-$+\sum\limits_{i=1}^{q}{\sum\limits_{j=1}^{q}{{{\cal L}_{Inter}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)}},$
+${{\cal L}_{stu}}={{\cal L}_{MRSTFT}}+\sum\limits_{k=1}^{q}{\sum\limits_{i=1}^{m}{\sum\limits_{j=1}^{n}{{{\cal L}_{Intra}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)}}}+\sum\limits_{i=1}^{q}{\sum\limits_{j=1}^{q}{{{\cal L}_{Inter}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)}},$ (11)
 where $q$ is the total number of correlated sets (set to 3 in this paper), $m$ is the number of layers contained in the student correlated set, and $n$ is the number of layers contained in the teacher correlated set.
-IV Experimental Settings
-IV-A Dataset Setup
-Input: Student and teacher correlated sets $\left\{{C_{s}^{k}}\right\}_{k=1}^{q}$ and $\left\{{C_{t}^{k}}\right\}_{k=1}^{q}$.
-Output: Intra-set and inter-set distillation loss ${{\cal L}_{Intra}}$ and ${{\cal L}_{Inter}}$.
-/* Iterate through all correlated sets. */
-1 for $k\leftarrow 1$ to $q$ do
-2       Sample intra-set student representations $\left({{\bf{F}}_{s}^{\rm{1}},{\bf{F}}_{s}^{2}\ldots,{\bf{F}}_{s}^{m}}\right)$ from $C_{s}^{k}$;
-3       Sample intra-set teacher representations $\left({{\bf{F}}_{t}^{\rm{1}},{\bf{F}}_{t}^{2}\ldots,{\bf{F}}_{t}^{n}}\right)$ from $C_{t}^{k}$;
-4       ${{\cal L}_{Intra}}=0$;
-       /* Iterate through all teacher-student layers in the current set. */
-5       for $i\leftarrow 1$ to $m$ do
-6             for $j\leftarrow 1$ to $n$ do
-7                   ${{\cal L}_{Intra}}+={\rm{TF\_Calibration}}\left({{\bf{F}}_{s}^{i},{\bf{F}}_{t}^{j}}\right)$;
-8                   if $i==m$ then // Ensure the uniqueness.
-9                         if $j==1$ then
-10                               ${\bf{R}}_{t}^{j}={\bf{F}}_{t}^{j}$;
-11
-12                        else
-13                               ${\bf{T}}_{t}^{j},{\bf{R}}_{t}^{j}={\rm{Fusion}}\left({{\bf{F}}_{t}^{j},{\bf{R}}_{t}^{j-1}}\right)$;
-14
-15                         end if
-16
-17                   end if
-18
-19             end for
-20            if $i==1$ then
-21                   ${\bf{R}}_{s}^{j}={\bf{F}}_{s}^{j}$;
-22
-23            else
-24                   ${\bf{T}}_{s}^{j},{\bf{R}}_{s}^{j}={\rm{Fusion}}\left({{\bf{F}}_{s}^{j},{\bf{R}}_{s}^{j-1}}\right)$;
-25
-26             end if
-27
-28       end for
-29      ${\bf{u}}_{s}^{k}={\bf{T}}_{s}^{m}$;
-       // The final output is chosen as the representative fused feature.
-30       ${\bf{u}}_{t}^{k}={\bf{T}}_{t}^{n}$;
-31
-32 end for
-[…]
-/* Iterate through fused features. */
-34 for $i\leftarrow 1$ to $q$ do
-35       for $j\leftarrow 1$ to $q$ do
-36             ${{\cal L}_{Inter}}+={\rm{TF\_Calibration}}\left({{\bf{u}}_{s}^{i},{\bf{u}}_{t}^{j}}\right)$;
-37
-38       end for
-39
-40 end for
-Algorithm 1 Intra-inter Set Distillation Processing
+4 Experimental settings
+4.1 Dataset setup
 The proposed distillation framework is trained and evaluated on two widely used public SE datasets: the single-channel deep noise suppression (DNS) challenge dataset [^1] The dataset is available at https://github.com/microsoft/DNS-Challenge.
@@
 Since the official blind test set is not fully open-sourced, we compare various algorithms on the 7-hour development test set provided by the challenge.
-IV-B Implementation Details
-TABLE I: Baseline model hyperparameter settings.
-Layer Teacher Student
+4.2 Implementation details
+Table 1: Baseline model hyperparameter settings.
+Layer
+Teacher Student
 Parameters Input Size Output Size Parameters Input Size Output Size
@@
 Conv2d (1, 3), (1, 2), 128 (128, $T$, $F/2$) (128, $T$, $F/4$) (1, 3), (1, 2), 64 (64, $T$, $F/2$) (64, $T$, $F/4$)
-Dilated Block (4$\times$×Conv2d) (2, 3), (1, 1), 128 $p$: [1, 2, 4, 8] (128, $T$, $F/4$) (128, $T$, $F/4$) $(2,3)$, $(1,1)$, 64 $p$: [1, 2, 4, 8] (64, $T$, $F/4$) (64, $T$, $F/4$)
-Frequency-Time Block (128 GRU×$\times$2)$\times$4 (128, $T$, $F/4$) (128, $T$, $F/4$) 64 GRU×$\times$2 (64, $T$, $F/4$) (64, $T$, $F/4$)
-Dilated Block (4$\times$×Conv2d) (2, 3), (1, 1), 128 $p$: [1, 2, 4, 8] (128, $T$, $F/4$) (128, $T$, $F/4$) $(2,3)$, $(1,1)$, 64 $p$: [1, 2, 4, 8] (64, $T$, $F/4$) (64, $T$, $F/4$)
+Dilated Block
+[…]
+(2, 3), (1, 1), 128
+$p$: [1, 2, 4, 8]
+(128, $T$, $F/4$)
+(128, $T$, $F/4$)
+$(2,3)$, $(1,1)$, 64
+$p$: [1, 2, 4, 8]
+(64, $T$, $F/4$)
+(64, $T$, $F/4$)
+[…]
+Dilated Block
+[…]
+(2, 3), (1, 1), 128
+$p$: [1, 2, 4, 8]
+(128, $T$, $F/4$)
+(128, $T$, $F/4$)
+$(2,3)$, $(1,1)$, 64
+$p$: [1, 2, 4, 8]
+(64, $T$, $F/4$)
+(64, $T$, $F/4$)
 DeConv2d (1, 3), (1, 2), 128 (128, $T$, $F/4$) (128, $T$, $F/2$) (1, 3), (1, 2), 64 (64, $T$, $F/4$) (64, $T$, $F/2$)
 DeConv2d (1, 3), (1, 2), 2 (128, $T$, $F/2$) (2, $T$, $F$) (1, 3), (1, 2), 2 (64, $T$, $F/2$) (2, $T$, $F$)
-TABLE II: Distillation module hyperparameter settings.
+Table 2: Distillation module hyperparameter settings.
 Similarity Embedding Mapping
-Layer Time Flow Frequency Flow
+Layer
+Time Flow Frequency Flow
 Parameters Input Size Output Size Parameters Input Size Output Size
-Linear $T*factor$ units ($B$, $T$, $T$) ($B$, $T$, $T*factor$) $B*factor$ units ($T$, $B$, $B$) ($T$, $B$, $B*factor$)
-ReLU - ($B$, $T$, $T*factor$) ($B$, $T$, $T*factor$) - ($T$, $B$, $B*factor$) ($T$, $B$, $B*factor$)
-Linear $T$ units ($B$, $T$, $T*factor$) ($B$, $T$, $T$) $B$ units ($T$, $B$, $B*factor$) ($T$, $B$, $B$)
-${l^{2}}$orm - ($B$, $T$, $T$) ($B$, $T$, $T$) - ($T$, $B$, $B$) ($T$, $B$, $B$)
+Linear TfactorT*factor units ($B$, $T$, $T$) ($B$, $T$, TfactorT*factor) BfactorB*factor units ($T$, $B$, $B$) ($T$, $B$, BfactorB*factor)
+ReLU - ($B$, $T$, TfactorT*factor) ($B$, $T$, TfactorT*factor) - ($T$, $B$, BfactorB*factor) ($T$, $B$, BfactorB*factor)
+Linear $T$ units ($B$, $T$, TfactorT*factor) ($B$, $T$, $T$) $B$ units ($T$, $B$, BfactorB*factor) ($T$, $B$, $B$)
+$l^{2}$}Norm - ($B$, $T$, $T$) ($B$, $T$, $T$) - ($T$, $B$, $B$) ($T$, $B$, $B$)
 Residual Fusion
-Layer Teacher Student
+Layer
+Teacher Student
 Parameters Input Size Output Size Parameters Input Size Output Size
@@
 Up Sampling - (128, $T$, $F_{t}^{j-1}$) (128, $T$, $F_{t}^{j}$) - (128, $T$, $F_{s}^{j-1}$) (64, $T$, $F_{s}^{j}$)
-Concatenate - 2$\times$(128, $T$, $F_{t}^{j}$) (256, $T$, $F_{t}^{j}$) - 2$\times$(64, $T$, $F_{s}^{j}$) (128, $T$, $F_{s}^{j}$)
+[…]
 Conv1d 1, 1, 2 (256, $T$, $F_{t}^{j}$) (2, $T$, $F_{t}^{j}$) 1, 1, 2 (128, $T$, $F_{s}^{j}$) (2, $T$, $F_{s}^{j}$)
-Sigmoid - (2, $T$, $F_{t}^{j}$) 2$\times$(1, $T$, $F_{t}^{j}$) - (2, $T$, $F_{s}^{j}$) 2$\times$(1, $T$, $F_{s}^{j}$)
+[…]
 Conv2d (3, 3), (1, 1), $C_{t}^{j}$ (128, $T$, $F_{t}^{j}$) ($C_{t}^{j}$, $T$, $F_{t}^{j}$) (3, 3), (1, 2), $C_{s}^{j}$ (64, $T$, $F_{s}^{j}$) ($C_{s}^{j}$, $T$, $F_{s}^{j}$)
@@
 For the single-channel DNS dataset, the input channels are set to 2, while for the multi-channel L3DAS23 dataset, the input channels are set to 16.
-The hyperparameters of the teacher and student models are shown in Table I, where $C$ represents the number of input channels, $T$ is the number of speech frames, and $F$ is the frequency dimension of speech features.
+The hyperparameters of the teacher and student models are shown in Table 1, where $C$ represents the number of input channels, $T$ is the number of speech frames, and $F$ is the frequency dimension of speech features.
 For the encoder and decoder parts, the number of channels in each convolutional layer of the teacher is twice that of the student.
-In Table I, $p$ represents the dilation factor, and both the teacher and student models use four types of dilation receptive fields.
+In Table 1, $p$ represents the dilation factor, and both the teacher and student models use four types of dilation receptive fields.
 For the intermediate part, the teacher model employs four F-T modules, while the student model uses only one F-T module, with the number of GRU hidden units also being half those of the teacher model’s.
-The hyperparameter settings for the trainable layers in the proposed distillation components are shown in Table II.
+The hyperparameter settings for the trainable layers in the proposed distillation components are shown in Table 2.
 The trainable components contain the computation of query and key embedding matrices in time-frequency calibration, as well as the generation process of recursively fused features for each correlated set.
 In the time-frequency calibration, the similarity embedding mappings are computed separately from two flow directions, with each flow containing two fully-connected layers for dimension transformation.
-The dimension transformation coefficient $factor$ in Table II is set to four in this paper.
+The dimension transformation coefficient factorfactor in Table 2 is set to four in this paper.
 Residual fusion integrates the current layer features with the recursive features inherited from the previous layer via trainable parameters.
-In Table II, $C_{t/s}^{j}$ represents the number of teacher-student channels in the current layer, $F_{t/s}^{j}$ denotes the feature dimension of teacher-student features in the current layer, and $F_{t/s}^{j-1}$ denotes the feature dimension in the previous layer.
+In Table 2, $C_{t/s}^{j}$ represents the number of teacher-student channels in the current layer, $F_{t/s}^{j}$ denotes the feature dimension of teacher-student features in the current layer, and $F_{t/s}^{j-1}$ denotes the feature dimension in the previous layer.
 The number of channels for recursive features of the teacher and student are set to 128 and 64, respectively.
@@
 The training configuration includes a learning rate of 0.0006 with the Adam optimizer, a batch size of 8, and a total training duration of 20 epochs.
-We provide source codes to ensure reproducibility [^3] The codes are available at https://github.com/JMCheng-SEU/I2S-TFCKD-SE..
-IV-C Comparative Methods
+4.3 Comparative methods
 We conduct multiple comparative experiments to comprehensively evaluate the proposed distillation framework for SE.
@@
 Spatial Net [23] primarily consists of interleaved narrow-band and cross-band blocks to respectively exploit narrow-band and cross-band spatial information.
-IV-D Evaluation Metrics
-In this paper, we use six objective speech quality evaluation metrics, including the perceptual evaluation of speech quality (PESQ) [46], short-time objective intelligibility (STOI) [47], scale-invariant signal-to-noise ratio (SI-SNR) [48], and three mean opinion ratings on quality scales [49]: signal distortion (CSIG), the intrusiveness of background noise (CBAK), and overall quality (COVL).
+4.4 Evaluation metrics
+In this paper, we use six objective speech quality evaluation metrics, including the perceptual evaluation of speech quality (PESQ) [^3] The evaluation codes are available at https://ecs.utdallas.edu/loizou/speech/software.htm., short-time objective intelligibility (STOI) [^4] The evaluation codes are available at http://www.ceestaal.nl/code/., scale-invariant signal-to-noise ratio (SI-SNR)3, and three mean opinion ratings on quality scales3: signal distortion (CSIG), the intrusiveness of background noise (CBAK), and overall quality (COVL).
 For all these metrics, higher scores indicate better performance.
-Additionally, on the L3DAS23 dataset, we use the official recommended word error rate (WER) computed via Wav2Vec [50] to evaluate SE effects for speech recognition purposes, where lower values indicate better performance.
-V Experimental Results and Discussion
-V-A Ablation Study
+Additionally, on the L3DAS23 dataset, we use the official recommended word error rate (WER) [^5] The evaluation codes are available at https://github.com/l3das/L3DAS22.
+to evaluate SE effects for speech recognition purposes, where lower values indicate better performance.
+5 Experimental results and discussion
+5.1 Ablation study
 We conduct ablation experiments for the distillation components on both the validation set and the multi-SNR test set to verify the effectiveness of the proposed modules step-by-step.
-First, we set the M1 method as the basic form of feature distillation, adopting the teacher-student layer-wise distillation framework of FitNet.
-Then, method M2 is built upon the cross-layer attention-calibration distillation framework proposed by SemCKD, leveraging multi-layer teacher feature information to guide the student model.
-Method M3 expands on M2’s framework by introducing the multi-layer matching distillation with time-frequency calibration proposed in Section III-C of this paper.
+First, we set the M1 (layer-wise KD) method as the basic form of feature distillation, adopting the teacher-student layer-wise distillation framework of FitNet [26].
+Then, M2 (intra-set matching with SemCKD [32]) is constructed on the basis of an intra-set cross-layer distillation framework, where multi-layer matching distillation was performed within each correlated set and calibrated by the cross-layer attention mechanism introduced in SemCKD [32].
+M3 (intra-set matching with TFCKD) further extended M2 by replacing the attention-based alignment with the proposed time–frequency cross-calibration described in Section 3.3 of this paper.
 Finally, the M4 method (i.
-e., the proposed I2S-TFCKD framework) adds correlated set residual fusion to M3, and performs intra-inter set multi-layer distillation.
+e., the proposed I2RF-TFCKD framework) performs time–frequency cross-calibrated multi-layer distillation both within and across correlated sets through recursive feature fusion, forming an intra–inter representation fusion distillation framework.
+The training trends of backbone loss and the speech quality evaluation metric PESQ on the validation set are illustrated in Fig.
+5(a) and Fig.
+5(b), respectively.
+Table 3 presents the comparison of objective speech metrics across different distillation strategies on the multi-SNR test set, and statistical significance analysis is conducted in the average metric comparison under three SNRs.
+We use the significance level $p$ to measure the statistical improvement of each distillation method compared to the student model.
 Figure 5: Training curve trends on the DNS validation set.
-Figure 5: Frame-level heatmap distribution of time-frequency alignment weights.
-V-A1 Analysis of Training Curves Trend
-Fig.
-4(a) shows the trend of each method’s backbone loss with the training epochs on the validation set.
-It can be observed that the feature distillation M1 enables the student’s backbone loss to decrease rapidly in the early training stages, but this initial advantage does not persist until model convergence.
-In contrast, method M2 achieves a lower converged backbone loss, indicating that cross-layer interactions provide additional knowledge to sustain the student’s learning.
-After introducing the time-frequency cross-calibration distillation loss in the M3 method, the variation of backbone loss with training epochs becomes more uniform and converges better, indicating that assigning multi-layer distillation weights based on time-frequency calibration leads to more stable knowledge transfer.
-On this basis, the M4 method introducing residual fusion distillation between correlated sets leads to lower convergence of the backbone loss.
-This demonstrates that intra-set and inter-set distillation complement each other, expanding the depth of knowledge perceived by the student model.
-Fig.
-4(b) shows the variation of the speech metric PESQ with training epochs.
-The general performance order of different distillation methods is M1<M2<M3<M4, consistent with that in Fig.
-4(a).
-V-A2 Heatmap Visualization Analysis of Time-frequency Calibration Weights
-We extract the trained time-frequency cross-calibration weights for a speech segment and compute the frame-level average energy along the feature dimension.
-Heatmap analysis is performed along the temporal and frequency domains, respectively, integrated with spectrograms of noisy and clean speech, as illustrated in Fig.
-5.
-It can be observed that the temporal calibration weights allocate higher weights to frame regions with speech activity, whereas the weights computed along the frequency domain show no obvious focusing bias.
-This indicates that the temporal calibration weights can capture contribution differences across different speech frames, whereas frequency branch focuses on analyzing distillation contributions between different layers within the current frame.
-V-A3 Analysis of Evaluation Metrics
-TABLE III: Speech evaluation metrics for ablation comparisons of distillation components under different SNRs.
-Methods Feature distillation Cross-layer distillation T-F calibration Intra-inter interactions Metric
+Table 3: Speech evaluation metrics for ablation comparisons of distillation components under different SNRs.
+Methods
+Feature
+distillation
+Cross-layer
+distillation
+T-F
+calibration
+Intra-inter
+interactions
+Metric
 PESQ CSIG CBAK COVL STOI SISNR
@@
 KD methods:
--M1 - - - 1.966 3.377 2.840 2.657 0.870 0.844
--M2 - - - 2.003 3.437 2.884 2.708 0.874 11.020
--M3 - - 2.057 3.500 2.894 2.767 0.880 11.107
--M4 - 2.106 3.567 2.950 2.830 0.886 11.253
+-M1 (layer-wise KD) - - - 1.966 3.377 2.840 2.657 0.870 0.844
+-M2 (intra-set matching with SemCKD [32]) - - - 2.003 3.437 2.884 2.708 0.874 11.020
+-M3 (intra-set matching with TFCKD) - - 2.057 3.500 2.894 2.767 0.880 11.107
+-M4 (I2RF-TFCKD) - 2.106 3.567 2.950 2.830 0.886 11.253
 0dB SNR
@@
 KD methods:
--M1 - - - 2.329 3.772 3.224 3.059 0.920 14.106
--M2 - - - 2.368 3.826 3.262 3.107 0.924 14.237
--M3 - - 2.426 3.904 3.274 3.174 0.928 14.270
--M4 - 2.491 3.954 3.333 3.238 0.931 14.461
+-M1 (layer-wise KD) - - - 2.329 3.772 3.224 3.059 0.920 14.106
+-M2 (intra-set matching with SemCKD [32]) - - - 2.368 3.826 3.262 3.107 0.924 14.237
+-M3 (intra-set matching with TFCKD) - - 2.426 3.904 3.274 3.174 0.928 14.270
+-M4 (I2RF-TFCKD) - 2.491 3.954 3.333 3.238 0.931 14.461
 5dB SNR
@@
 KD methods:
--M1 - - - 2.700 4.147 3.591 3.452 0.952 17.041
--M2 - - - 2.724 4.180 3.617 3.483 0.953 17.149
--M3 - - 2.788 4.244 3.633 3.546 0.956 17.157
--M4 - 2.850 4.285 3.686 3.606 0.958 17.356
+-M1 (Layer-wise KD) - - - 2.700 4.147 3.591 3.452 0.952 17.041
+-M2 (Intra-set matching with SemCKD [32]) - - - 2.724 4.180 3.617 3.483 0.953 17.149
+-M3 (Intra-set matching with TFCKD) - - 2.788 4.244 3.633 3.546 0.956 17.157
+-M4 (I2RF-TFCKD) - 2.850 4.285 3.686 3.606 0.958 17.356
 Average SNR
@@
 KD methods:
--M1 - - - 2.332 ($\ast$)1 3.765 ($\ast$) 3.218 ($\ast$) 3.056 ($\ast$) 0.914 ($\ast$) 13.997 ($\ast$)
-[…]
-[…]
-[…]
+-M1 (Layer-wise KD) - - - 2.332 ($\ast$)1 3.765 ($\ast$) 3.218 ($\ast$) 3.056 ($\ast$) 0.914 ($\ast$) 13.997 ($\ast$)
+-M2 Intra-set matching with SemCKD [32] - - - 2.365 ($\ast$) 3.815 ($\ast\,\ast$) 3.254 ($\ast$) 3.099 ($\ast$) 0.919 ($\ast\,\ast$) 14.135 ($\ast$)
+-M3 (Intra-set matching with TFCKD) - - 2.424 ($\ast\,\ast$) 3.883 ($\ast\ast\ast$) 3.267 ($\ast\,\ast$) 3.162 ($\ast\ast\ast$) 0.921 ($\ast\ast\ast$) 14.178 ($\ast$)
+-M4 (I2RF-TFCKD) - 2.482 ($\ast\ast\ast$) 3.935 ($\ast\ast\ast$) 3.323 ($\ast\ast\ast$) 3.225 ($\ast\ast\ast$) 0.925 ($\ast\ast\ast$) 14.357 ($\ast$)
   • 1
-[…]
-Table III presents the comparison of objective speech metrics across different distillation strategies on the multi-SNR test set, and statistical significance analysis is conducted in the average metric comparison under three SNRs.
-We use the significance level $p$ to measure the statistical improvement of each distillation method compared to the student model.
-From the results, all distillation methods improve the student’s objective metrics across the three SNR levels, narrowing the gap between the teacher and student.
-Among them, M1 that performs layer-wise distillation, underperforms other methods, showing no significant improvement ($p>0.1$) over the student model in any of the six metrics.
-In contrast, M2 that extends single-layer distillation to attention-based multi-layer distillation, outperforms M1 across all metrics and achieves secondarily significant improvements ($0.10.05<p<$ 0.1) in CSIG and STOI compared to the student model.
-This indicates that cross-layer interactions provide additional representational knowledge.
-Building on this, M3 that introduces time-frequency cross-calibration, achieves further metric advantages, with statistically significant improvements ($p<0.05$) in CSIG, COVL, and STOI compared to the student model.
-This demonstrates that cross-calibration captures discriminative information across frequency bands and temporal segments at different layers, enabling more refined teacher-student layer matching.
-M4 that introduces residual fusion for inter-set distillation, achieves the best metric results among all strategies, with statistically significant improvements ($p<0.05$) in five metrics under average SNR compared to the student, except for SISNR.
-This indicates that collaborative distillation within and between correlated sets effectively promotes global knowledge circulation, significantly improving the enhancement effect of the student model.
-V-B Comparative Experiments on the DNS Benchmark
-TABLE IV: Comparison of objective speech evaluation metrics between distilled and undistilled models on the DNS test set.
-Methods Casual & NFI Param. (M) FLOPs (G/s) PESQ STOI SISNR (dB)
-Noisy - - - 1.58 0.915 9.07
-RNNoise [34$\dagger$1 0.1 0.04 1.97 - -
-NSNet [35$\dagger$ 1.3 0.13 2.15 0.945 15.61
-DTLN [36$\dagger$ 1.0 0.25 - 0.948 16.34
-DCCRN [37$\dagger$ 3.7 14.37 2.80 0.964 16.38
-FullSubNet+ [38$\dagger$ 8.7 31.12 2.98 0.967 18.34
-FRCRN [6$\dagger$ 6.9 12.30 3.23 0.977 19.78
-CTS-Net [39$\dagger$ 4.4 5.57 2.94 0.967 17.99
-PrimeK-Net [40$\ddagger$ 1.4 44.64 3.34 0.978 19.23
-DPDCRN-T 3.5 13.71 3.16 0.973 17.47
-DPDCRN-S 0.6 2.44 2.81 0.962 17.05
-KD methods
--MSE [32$\ddagger$ 0.6 2.44 2.82 0.962 16.60
--CLSKD [26$\dagger$ 0.6 2.44 2.84 0.963 17.08
--UCLFWPKD [16$\dagger$ 0.6 2.44 2.94 0.966 17.16
--ABC-KD [14$\ddagger$ 0.6 2.44 2.90 0.964 17.14
--Two-Step KD [15$\ddagger$ 0.6 2.44 2.93 0.965 17.05
--M1 (FitNet [33$\ddagger$) 0.6 2.44 2.83 0.962 17.22
--M2 (SemCKD [30$\ddagger$) 0.6 2.44 2.89 0.965 17.36
--M3 (TFCKD) 0.6 2.44 2.97 0.967 17.58
--M4 (I2S-TFCKD) 0.6 2.44 3.03 0.968 17.74
-  • 1
-    The model results labeled with symbol $\dagger$ are taken directly from the original paper, while the symbol $\ddagger$ denotes the results reproduced on the datasets of this paper according to the original paper’s settings.
+    The significance analysis [7] of each distillation method and the baseline DPDCRN-S under average SNR is presented by the symbol $\ast$, and the significance level $p$ is: $\ast$: $p>0.1$, $\ast\,\ast$: $0.05<p<0.1$, $\ast\ast\ast$ : $p<0.05$.
+5.1.1 The impact of intra-set cross-layer distillation
+Compared with layer-wise KD M1, M2 further introduces an intra-set cross-layer distillation mechanism that enables the student to absorb multi-level hierarchical knowledge from the teacher.
+As shown in Fig.
+5(a), M1 enables the student’s backbone loss to decrease rapidly in the early training stages, but this initial advantage does not persist until model convergence.
+In contrast, M2 achieves a lower converged backbone loss, indicating that cross-layer interactions provide additional knowledge to sustain the student’s learning.
+In the PESQ curves of Fig.
+5(b), M2 consistently outperforms M1, which further validates this observation.
+In terms of the metric comparisons on the multi-SNR test set in Table 3, M1 is inferior to other methods, showing no significant improvement ($p>0.1$) over the student model in any of the six metrics.
+By contrast, M2 that extends distillation to multi-layer teacher–student interactions, outperforms M1 across all metrics and achieves secondarily significant improvements ($0.05<p<0.1$) in CSIG and STOI compared to the student model.
+These results demonstrate that the additional representational information provided by cross-layer interactions offers constructive guidance to the student in both speech quality and noise suppression.
+5.1.2 The impact of time–frequency cross-calibration
+Building upon the cross-layer KD framework of M2, M3 introduces the proposed time–frequency cross-calibration mechanism, which adjusts the distillation weights across layers tailored to the characteristics of speech signals.
+As shown in Fig.
+5(a), the backbone loss of M3 exhibits a more uniform trend and achieves better convergence compared with M2, indicating that assigning multi-layer distillation weights based on time-frequency calibration leads to more stable knowledge transfer.
+Similarly, M3 also achieves a better performance ceiling on the PESQ curve of the validation set.
+Regarding the objective metrics on multi-SNR test set, M3 achieves statistically significant improvements over the student model on CSIG, COVL, and STOI ($p<0.05$), and consistently outperforms both M1 and M2 across all three SNR conditions.
+These observations demonstrate that cross-calibration captures discriminative information across frequency bands and temporal segments at different layers, enabling more refined teacher-student layer matching.
+5.1.3 The impact of the intra-inter representation fusion framework
+Distinct from the previous three methods, M4 further introduces a global knowledge circulation mechanism across both intra-set and inter-set levels.
+This framework not only enables multi-layer teacher–student feature interactions within a single correlated set, but also facilitates cross-set deep fusion of teacher and student representations.
+As shown in the training curves of Fig.
+5, M4 achieves the lowest convergence point of backbone loss among all schemes, while its PESQ curve consistently remains at the highest level, demonstrating stable and sustained performance gains.
+This indicates that intra-set and inter-set distillation complement each other, expanding the depth of knowledge perceived by the student model.
+Considering the objective metrics comparisons in Table 3, M4 achieves statistically significant improvements ($p<0.05$) over the student model on five metrics under average SNR, except for SISNR, and shows clear advantages over the other methods across all SNR levels.
+In summary, the proposed M4 method maximizes the transferability of the teacher by ensuring the student’s absorption of intra-set fine-grained knowledge while simultaneously broadening its understanding of the teacher’s inter-set global information.
+5.1.4 Heatmap visualization analysis of time-frequency calibration weights
+We extract the trained time-frequency cross-calibration weights for a speech segment and compute the frame-level average energy along the feature dimension.
+Heatmap analysis is performed along the temporal and frequency domains, respectively, integrated with spectrograms of noisy and clean speech, as illustrated in Fig.
+6.
+It can be observed that the temporal calibration weights allocate higher weights to frame regions with speech activity, whereas the weights computed along the frequency domain show no obvious focusing bias.
+This indicates that the temporal calibration weights can capture contribution differences across different speech frames, whereas frequency branch focuses on analyzing distillation contributions between different layers within the current frame.
+Figure 6: Frame-level heatmap distribution of time-frequency alignment weights.
+5.2 Comparative experiments on the DNS benchmark
 Fair comparisons of various algorithms are conducted on the DNS non-reverb test set.
-Table IV presents the objective speech quality evaluation metrics of various SE methods.
+Table 4 presents the objective speech quality evaluation metrics of various SE methods.
 We provide details on the causality of each model and whether it uses future frame information (Casual and No Future Information, Casual & NFI).
@@
 Additionally, the model parameter (Param.) count and floating-point operations (FLOPs) per second are provided to evaluate deployment feasibility on edge devices.
-Among the SE models in Table IV, NSNet, RNNoise, and DTLN are low-complexity real-time SE algorithms primarily based on LSTM networks.
+Among the SE models in Table 4, NSNet, RNNoise, and DTLN are low-complexity real-time SE algorithms primarily based on LSTM networks.
 These models are fully causal and do not utilize future information.
@@
 Our distillation baseline DPDCRN, composed of causal convolutions and GRU layers, is strictly causal and operates without any future frame information.
-V-B1 Analysis of Distillation Frameworks
-Figure 6: The scatter-bubble distribution of PESQ, FLOPs, and model parameters for SE models on the DNS test set.
-We compare the effects of different distillation frameworks applied to the base model DPDCRN in Table IV.
+5.2.1 Analysis of distillation frameworks
+Figure 7: The scatter-bubble distribution of three metrics (PESQ, STOI, and SISNR) with model parameters and FLOPs for SE models on the DNS test set.
+We compare the effects of different distillation frameworks applied to the base model DPDCRN in Table 4.
 Among them, MSE that directly narrows the output distance, only achieves slightly higher results than the student model on PESQ and STOI metrics and even falls below the student on the SISNR metric.
@@
 This highlights that computing multi-layer calibration weights along both the temporal and spectral domains enables more effective knowledge allocation.
-Extending M3 to the intra-inter set distillation framework (M4, I2S-TFCKD) achieves the best performance among all distillation methods, demonstrating that knowledge circulation within and between correlated sets further facilitate teacher-student knowledge transfer.
-V-B2 Comparison with State-of-the-art SE Methods
-Table IV provides the metric performances of current SOTA SE algorithms on the DNS test set.
+Extending M3 to the intra-inter set distillation framework (M4, I2RF-TFCKD) achieves the best performance among all distillation methods, demonstrating that knowledge circulation within and between correlated sets further facilitate teacher-student knowledge transfer.
+5.2.2 Comparison with state-of-the-art SE methods
+Table 4: Comparison of objective speech evaluation metrics between distilled and undistilled models on the DNS test set.
+Methods Casual & NFI Param. (M) FLOPs (G/s) PESQ STOI SISNR (dB)
+Noisy - - - 1.58 0.915 9.07
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+DPDCRN-T 3.5 13.71 3.16 0.973 17.47
+DPDCRN-S 0.6 2.44 2.81 0.962 17.05
+KD methods
+[…]
+[…]
+[…]
+[…]
+[…]
+-M1 (Layer-wise KD) 0.6 2.44 2.83 0.962 17.22
+[…]
+-M3 (Intra-set matching with TFCKD) 0.6 2.44 2.97 0.967 17.58
+-M4 (I2RF-TFCKD) 0.6 2.44 3.03 0.968 17.74
+  • 1
+    The model results labeled with symbol $\dagger$ are taken directly from the original paper, while the symbol $\ddagger$ denotes the results reproduced on the datasets of this paper according to the original paper’s settings.
+Table 4 provides the metric performances of current SOTA SE algorithms on the DNS test set.
 Additionally, Fig.
-6 shows the scatter-bubble distribution of PESQ with model parameters and FLOPs.
+7 shows the scatter-bubble distribution of three metrics (PESQ, STOI, and SISNR) with model parameters and FLOPs.
 It can be observed that current real-time SE algorithms such as RNNoise, NSNet, and DTLN, despite having extremely low parameters and FLOPs, still exhibit a large performance gap compared to high-complexity SE models.
@@
 The recently proposed SE model PrimeK-Net achieves substantial parameter reduction through targeted architectural design, but its 44.64G FLOPs still impose enormous deployment pressure.
-This paper applies intra-inter set distillation with time-frequency calibration (I2S-TFCKD) to the student model DPDCRN-S, attaining competitive results across three metrics with low parameter count (0.6M) and computational complexity (2.4G FLOPs).
-V-C Comparative Experiments on the L3DAS23 Benchmark
-Figure 7: Boxplot of PESQ and the T1 Metric for different distillation strategies on the L3DAS23 validation and development sets.
-V-C1 Boxplot Analysis of Distillation Strategies
+This paper applies intra-inter set distillation with time-frequency calibration (I2RF-TFCKD) to the student model DPDCRN-S, attaining competitive results across three metrics with low parameter count (0.6M) and computational complexity (2.4G FLOPs).
+5.3 Comparative experiments on the L3DAS23 benchmark
+5.3.1 Boxplot analysis of distillation strategies
 Fig.
@@
 From the perspective of the mean value of the metrics, MSE still remains inferior to feature-based distillation strategies, indicating that the utilization of intermediate representation information is also crucial for multi-channel SE distillation.
-Among all distillation algorithms, the proposed I2S-TFCKD method achieves the highest mean values across both metrics.
+Among all distillation algorithms, the proposed I2RF-TFCKD method achieves the highest mean values across both metrics.
 Compared with the second-highest performing UCLFWPKD, the proposed framework has advantages in both the average metric and the concentration degree of high values, further validating the positive contributions of intra-inter set knowledge flow and time-frequency cross-calibration to the overall distillation effect.
-V-C2 Metric comparison Across Models
-Table V presents objective metric results on the L3DAS23 development set, comparing multi-channel SE models with and without KD.
+5.3.2 Metric comparison across models
+Table 5: Comparison of objective speech evaluation metrics between distilled and undistilled models on the development set in the SE track of L3DAS23.
+[…]
+Noisy1 - - - 1.166 0.474 0.575 0.550
+[…]
+[…]
+[…]
+[…]
+[…]
+DPDCRN-T 3.5 13.71 2.036 0.129 0.885 0.878
+DPDCRN-S 0.6 2.44 1.711 0.185 0.850 0.832
+KD methods
+[…]
+[…]
+[…]
+[…]
+[…]
+-M1 (Layer-wise KD) 0.6 2.44 1.774 0.179 0.850 0.836
+[…]
+-M3 (Intra-set matching with TFCKD) 0.6 2.44 1.914 0.152 0.866 0.857
+-M4 (I2RF-TFCKD) 0.6 2.44 1.929 0.150 0.870 0.860
+  • 1
+    Note: Noisy data is taken from the first channel of microphone A.
+Figure 8: Boxplot of PESQ and the T1 Metric for different distillation strategies on the L3DAS23 validation and development sets.
+Table 5 presents objective metric results on the L3DAS23 development set, comparing multi-channel SE models with and without KD.
 The base model DPDCRN achieved first place on the blind test set of the L3DAS23 challenge and is causally adjusted in this paper.
@@
 Application of various distillation strategies to the student results in varying degrees of improvement across metrics.
-The proposed I2S-TFCKD method achieves optimal distillation effect, with indicator gains of 0.218 in PESQ, 0.035 in WER, 0.02 in STOI, and 0.028 in the T1 Metric compared to the undistilled student.
+The proposed I2RF-TFCKD method achieves optimal distillation effect, with indicator gains of 0.218 in PESQ, 0.035 in WER, 0.02 in STOI, and 0.028 in the T1 Metric compared to the undistilled student.
 Overall, the distilled student model maintains causal inference with low computational overhead while achieving competitive performance compared to SOTA SE methods in enhancement results.
-TABLE V: Comparison of objective speech evaluation metrics between distilled and undistilled models on the development set in the SE track of L3DAS23.
-[…]
-Noisy1 - - - 1.166 0.474 0.575 0.550
-Neural Beamforming [41$\dagger$ 5.5 32.15 - 0.569 0.684 0.557
-EaBNet [42$\dagger$ 2.8 7.38 - 0.549 0.724 0.587
-CCA Speech [43$\dagger$ - - - 0.293 0.836 0.771
-DeFT-AN [44$\ddagger$ 2.7 37.99 1.976 0.137 0.866 0.864
-Spatial Net [45$\ddagger$ 1.6 46.30 2.264 0.133 0.910 0.889
-DPDCRN-T 3.5 13.71 2.036 0.129 0.885 0.878
-DPDCRN-S 0.6 2.44 1.711 0.185 0.850 0.832
-KD methods
--MSE [32$\ddagger$ 0.6 2.44 1.759 0.181 0.847 0.833
--CLSKD [26$\dagger$ 0.6 2.44 1.768 0.178 0.851 0.837
--UCLFWPKD [16$\dagger$ 0.6 2.44 1.852 0.162 0.861 0.850
--ABC-KD [14$\ddagger$ 0.6 2.44 1.795 0.169 0.856 0.843
--Two-Step KD [15$\ddagger$ 0.6 2.44 1.832 0.169 0.859 0.845
--M1 (FitNet [33$\ddagger$) 0.6 2.44 1.774 0.179 0.850 0.836
--M2 (SemCKD [30$\ddagger$) 0.6 2.44 1.851 0.161 0.859 0.849
--M3 (TFCKD) 0.6 2.44 1.914 0.152 0.866 0.857
--M4 (I2S-TFCKD) 0.6 2.44 1.929 0.150 0.870 0.860
-  • 1
-    Note: Noisy data is taken from the first channel of microphone A.
-VI Conclusion
-In this paper, we proposed an intra-inter set distillation framework with time-frequency calibration (I2S-TFCKD) for SE model compression.
-Specifically, I2S-TFCKD partitions intermediate layers of SE models into multiple correlated sets.
+6 Conclusion
+In this paper, we proposed an intra-inter set distillation framework with time-frequency calibration (I2RF-TFCKD) for SE model compression.
+Specifically, I2RF-TFCKD partitions intermediate layers of SE models into multiple correlated sets.
 Within each set, intra-set cross-layer knowledge transfer is performed, and representative features are generated through residual fusion.
@@
 Experimental results on two public datasets (DNS and L3DAS23) demonstrated that the proposed distillation framework stably and effectively improves the enhancement performance of the low-complexity student model, narrows the gap with the teacher and outperforms other distillation strategies across various metrics.
-Future research will focus on multi-teacher distillation for SE models, enabling an elite student to handle various acoustic scenarios through knowledge integration from multiple teachers.
+Despite the progress achieved by our distillation method in the compression of SE models, there are still some limitations that need to be addressed.
+For example, the performance of our method on tasks such as speech separation and echo cancellation requires further validation.
+In addition, it remains unclear whether our method can be extended to multi-teacher distillation scenarios.
+In future work, we will continue to explore the application of the proposed method to more diverse speech signal processing tasks, and further investigate multi-teacher distillation for speech models.
+7 Acknowledgments
+This work was supported in part by the Natural Science Foundation of the Jiangsu Higher Education Institutions of China under Grant No.
+25KJB510010, and the National Natural Science Foundation of China under Grant No.
+62472227 and No.
+61871213.